### v1

# Scaling of poroelastic coarsening and elastic arrest in crosslinked gels

Abstract
Recent experiments on crosslinked gels quenched from solvent-rich to solvent-poor conditions, show solvent-rich domains surrounded by gel-rich regions that coarsen followed by kinetic arrest at micron scales that persists for hours before full drainage to macroscopic equilibrium occurs on day timescales. Motivated by this, we present a general model for both coarsening and eventual arrest of pressure-driven coarsening in gels with solvent flow driven by interfacial tension of the solvent-rich domains and the gel. In gels in their viscoelastic time regime, this capillary force is dissipated by solvent transport through crosslinked polymers where the polymers develop transient elastic stresses due to solvent flow. At longer times, the long-ranged, elastic response of the gel provides a deterministic force which balances the capillary force thus arresting pressure-driven coarsening at a stiffness-dependent length. For a melt-like, polymer-rich gel, the coarsening length $\sim t^{1/4}$ where $t$ is the time, with an amplitude $\sim G^{-1/2}$, where $G$ is the shear modulus. The arrest length in this regime also scales $\sim G^{-1/2}$. For low polymer fractions where the dominant length scale is the mesh size, the coarsening length can scale as $t^{1/3}$ varying with $G^{-1/3}$ (as does the arrest length). Our predictions for the arrest-length scaling with $\sim G^{-1/2}$ for the melt-like gel are consistent with the measurements.
I Introduction
Recent experiments [9] on phase separation of PDMS cross-linked polymers swollen by an organic solvent (a polymer gel system) have shown results that are strikingly different from our simple understanding of phase separation in simpler, two-component systems. The cross-linked polymer PDMS gel is initially prepared at a higher temperature with a certain fraction of solvent in equilibrium. The temperature is then lowered, and the equilibrium solvent fraction within the gel decreases. Once the gel comes to equilibrium—which only occurs macroscopically after a day or two—the gel and the solvent exhibit phase separation.
At its lower temperature state, the gel initially contains excess solvent due to its initial high-temperature preparation. After an initial period of spinodal decomposition and nucleation, mesoscale domains of solvent-rich regions form in the gel, surrounded by gel that is polymer-rich and has less solvent than its initial, high temperature state. The domain structure coarsens (with small domains shrinking and large ones growing), and this process is reminiscent of Lifshitz-Slyozov-Wagner (LSW) [15, 27, 3], coarsening in the phase separation of two-component systems. Strikingly, in these experiments the coarsening is observed to undergo long-lived kinetic arrest at submicron or micron length scales, depending on the rigidity of the cross-linked gel. This mesoscale structure remains stable for several hours until, on the scale of days, all excess solvent eventually drains from the solvent-rich domains into a reservoir. The gel then becomes homogeneous at its low-temperature state with a lower solvent fraction than its prepared state.
The origin of this metastable, elastic arrest has interested several scientific groups. The main experimental observation [9] is that the long-time mesoscale domain size depends on $1/\sqrt{G}$, where $G$ is the shear modulus of the gel, as measured in six gels with $10\text{kPa}\lesssim G\lesssim 400\text{kPa}$. The coarsening kinetics were treated numerically using a Cahn-Hilliard type model (see SI of [9] ), but no experimental quantification was presented. However, the paper states that for the more rigid gels, no coarsening was observed on the time scale of the measurements, while for the three softest gels, coarsening of the domains was observed.
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects. They do not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach and leads to the elastic arrest prediction of scaling with $1/G^{1/2}$, in agreement with the experiments [9]. A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus (see Appendix for further discussion). A more recent theory of phase separation in elastic networks [20], based on coupled chemical-potential gradients and elastic stress, concluded that local elasticity alone cannot arrest coarsening, and that non-local elastic interactions are required. By contrast, our approach focuses explicitly on solvent flow through the gel coupled to its poroelastic response. By treating solvent permeation (Darcy flow) and elastic back-reaction directly—rather than incorporating elasticity solely through constitutive stress–energy relations—we demonstrate that elastic arrest of coarsening can occur even for linearly elastic gels, as explained below.
Other theories [21, 16, 10] proposed to explain these phenomena using equilibrium models, unrelated to coarsening and poroelastic flow. However, these studies do not account for: (a) Coarsening dynamics and its relation to solvent flow in the porous and deformable gel. (b) The fact that the system ultimately reaches equilibrium as a homogeneous gel which suggests slow drainage of solvent into the reservoir. They, as well as the models of [17, 5, 20] do not relate to the Darcy-like solvent flow in the gel, driven by the capillary pressure of the solvent-rich domain.
Here, we present a unified scaling theory of both domain coarsening and elastic arrest in crosslinked gels which exhibit crosslink-controlled elastic behavior at long times and a viscoelastic response activated by solvent flow through the network at intermediate times. Darcy flow [26, 2, 13] of the solvent in the gel occurs at time scales for which poroelastic dissipation, acting within the viscoelastic regime of the polymer network, controls the rate of coarsening, while the onset of long-time elastic response governs kinetic arrest [8, 17, 26]. When the characteristic coarsening time becomes comparable to the relaxation time separating flowing, viscoelastic regime from the solid-like elastic regime, further capillary-driven domain evolution is arrested by the development of long-range elastic stress in the gel. At the level of the Laplace force balance at the solvent-rich domain boundary, this elastic stress cancels the capillary driving force and halts further solvent flow. The resulting arrested domain size is therefore not set directly by microscopic gel length scales, but by a kinetic criterion and is amplified by a magnification factor arising from the large separation between the poroelastic time governing solvent transport and the much longer time required for the onset of elastic response. During this extended viscoelastic window, solvent can be transferred over progressively larger distances, converting nanoscale gel properties into mesoscale domain sizes at arrest.
The complexity of the problem is simplified by our scaling approach focused on the dimensionally correct dependence of the coarsening and elastic arrest in terms of the gel properties, without the fine distinctions of numerical factors or dimensionless factors of order unity [6]. Thus, instead of equalities, we use the convention $\sim$ to mean “scales as”. The theory predicts a coarsening exponent of $1/4$ for slow, poroelastic diffusion and the arrest size of the domains scales as $1/\sqrt{G}$; the latter agrees with the experiments. Finally, we note here that while the explanations below consider the case of shrinking, solvent-rich domains, our model and its scaling predictions apply to both shrinking and growing domains, which are of course related by conservation of solvent. When we discuss solvent flowing out of shrinking solvent-rich domains, the same physics applies to solvent being drawn into growing domains.
II Methods
II.1 LSW coarsening
For contrast with the coarsening of solvent-rich domains in gels, we briefly review the scaling arguments that lead to domains that grow or shrink as $t^{1/3}$[3]. We consider a binary (AB) liquid that has been quenched from its mixed phase to a temperature where the equilibrium state is one of macroscopic phase separation. In general, each of those phases contains both A and B and for simplicity, we denote the phase by its majority component (e.g., B-rich is termed B). At early times, the system undergoes spinodal decomposition followed by nucleation of domains of B, for example, of one liquid in the background of the other [3, 7].
These domains coarsen [3] and the kinetics of coarsening begins with the conservation relation that equates the number of molecules $N_{B}$ leaving (for example) a shrinking B domain (and eventually being drawn into a growing B domain), to their flux $J_{BA}$ at the interface of the B domain with the background of A: $dN_{B}/dt=-J_{BA}S$, where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. For scaling purposes, we can ignore the vectors, keeping in mind that the flux and velocity of the B-molecules exiting the B-rich domain are in the direction normal to the domain boundary. Here $S$ is the area of the shrinking domain which, for simplicity, we have taken to be spherical and henceforth we denote the time derivatives by a dot. $\dot{N}_{B}$ is related to the composition difference $\Delta c$ between the two phases and the change in the domain interface position and the domain area. The flux is the number of B molecules that exit the B-rich domain, per unit area per unit time and is written in terms of the B-molecule velocity as: $J_{BA}=c_{B}\ v$, where $c_{B}$ is the concentration (number density) of B molecules in the B-rich domain. For a shrinking domain, $\dot{N}_{B}<0$ implies that $J_{BA}>0$; the velocity is directed away from the domain. In scaling, we write: $\dot{N}_{B}\sim-\Delta c\,S\,\dot{\lambda}$ where $\lambda$ is the size of the $B$ domain. For a more detailed discussion of this and the following see Appendix C.
[…]
Using the conservation equation, we now write a scaling relation: $\Delta c\ \dot{\lambda}\sim-D\ \nabla\sigma/\lambda$ where $\Delta c$ is the concentration difference of the B molecules in the B-rich phase and the A-rich phase. $\nabla$ now is the extent of the chemical potential gradient in the direction normal to the surface of the B domain, evaluated at the domain interface. LSW scaling for the domain growth is obtained by approximating the gradient term by $\sigma/\lambda^{2}$ which originates in the solution of the diffusion equation for the B molecules in the background of A, at the domain surface [3]. This equation in the fast-diffusion limit (applicable to molecular diffusion in a liquid relative to the coarsening times), results in an equation for the spatial profile of the B molecule concentration in the A-rich domain where $\lambda$ determines the scale of the gradient. Below, we generalize the approximation for the gradient to include diffusion.
The result for a shrinking domain gives $\dot{\lambda}\sim-(D/\Delta c)\,\sigma/\lambda^{2}$. For scaling, we can take $\Delta c\sim c_{B}$ (not valid near a critical point) which gives the LSW scaling, $\lambda\sim\left(\ (D\ \sigma/c_{B})\ t\right)^{1/3}$, where we do not distinguish shrinking from growing vis a vis the scaling; the two are simply related by conservation of the B molecules. As discussed below, the equations for the chemical potential gradients can be transformed into equations for the osmotic pressure gradients of the B molecules in the A-rich phase.
II.2 Generalizing LSW to coarsening of solvent-rich domains in gels:
In the following, the gel properties such as interfacial tension with a solvent-rich domain, elastic modulus $G$, mesh size $\xi$ are envisioned as those that characterize the equilibrium gel. We again begin with the conservation relation, $\dot{N}_{B}\sim-\Delta c_{s}\,S\,\dot{\lambda}=-JS$ where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary. $\Delta c_{s}$ is the difference between the solvent concentration in the solvent-rich domain and the gel. The flux is given by $\vec{J}=c_{s}\vec{v}$ where $c_{s}$ solvent concentration in the solvent-rich domain which, due to capillary pressure, ejects solvent molecules into the gel with a $\vec{v}$. For a more detailed discussion of this and the following see Appendix C.
However, instead of diffusing in the polymer-rich phase, the velocity is governed by solvent flow in the deformable gel considered as a porous medium where the Darcy relation [6, 1, 2, 24] that connects the velocity to the pressure gradient: $\vec{v}=(k_{0}/\eta_{D})\nabla p$ where $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity appropriate to the solvent flow impeded by its friction with the polymer network. Because this transport is controlled by the hydrodynamic permeability of the solvent rather than by polymer rearrangements, the network topology is effectively fixed on the Darcy-flow timescale: polymer configurations do not relax rapidly enough to eliminate the stress induced by solvent flow. Physically, this reflects a separation of timescales between the relatively fast permeation of small solvent molecules through the gel and the much slower relaxation of stress carried by long-chain polymers. In the Results, we discuss specific choices for these parameters.
The pressure gradient entering Darcy’s law is the local gradient at the interface of the solvent-rich domain and the gel, whose magnitude is set by how the interfacial pressure perturbation spreads into the gel. In incompressible systems pressure gradients are proportional to chemical potential gradients of the solvent molecules in the gel, by the thermodynamic relation, $\nabla\mu=\nabla p/c_{B}$; see Eq. 1.1.14 of Ref. [19] and Appendix C. The chemical potential at the solvent-rich domain surface is again given by capillarity, $\delta\mu=\sigma/(\lambda c_{s})$ and the osmotic pressure jump at the interface (which acts as the mechanical pressure jump) is given by the Laplace law as $\delta p=c_{s}\delta\mu=\sigma/\lambda$. Finally, the conservation law and the Darcy relation used in the flux give the dynamical scaling relation for a shrinking domain:
$\dot{\lambda}\sim-\frac{k_{0}\ c_{s}}{\eta_{D}}\nabla p\sim-\frac{k_{0}\ \sigma}{\eta_{D}\ \lambda\ \ell_{p}}$ (1)
where $\ell_{p}$ is the scale of the pressure gradient, evaluated at the interface of the solvent-rich domain and the gel, and where we have set $c_{s}/\Delta c_{s}$ to be a dimensionless constant of order unity for scaling purposes.
The solvent pressure in the gel decreases from its maximal value at the boundary of the shrinking domain via stress diffusion, known as poroelasticity [24, 4]. That both the stress and pressure propagate diffusively can be understood from the low-Reynolds number force balance in a gel which implies that force balance implies that pressure gradients are balanced by stress gradients. At relatively early times, when the poroelastic diffusion length $\ell_{p}$ is smaller than the domain size $\lambda$, the pressure gradient at the domain boundary can be estimated as $\nabla p\sim\delta p/\ell_{p}$; this approximation captures the transient buildup of the interfacial gradient before it reaches its long-time, saturated value. As defined by poroelastic diffusion [26]:
$\ell_{p}(t)=\sqrt{D_{p}\,t},$ (2)
where $D_{p}\sim k_{0}\,G/\eta_{ps}$ is the poroelastic diffusion constant, with $k_{0}$ the permeability of the gel, $G$ its elastic modulus, and $\eta_{ps}$ the viscosity associated with the viscoelastic response of the coupled polymer–solvent system activated by solvent flow at intermediate times. This description applies only until $\ell_{p}$ reaches the domain size $\lambda(t)$. Beyond this point, the pressure field has equilibrated across the entire solvent-rich domain, and extending the gradient over larger length scales does not increase the capillary driving force acting on that domain. Mathematically, this is because the finite domain size, together with the Laplace pressure boundary condition at the domain interface, fix both the magnitude and the spatial extent of the pressure gradients. Those are substantial only for distances from the domain smaller or equal to $\lambda$. As shown in more detail in Appendix C, beyond this point, the solution of the poroelastic diffusion equation has evolved so that the pressure gradient evaluated at the domain interface no longer changes with time, and further spreading of the pressure field to larger distances does not increase the capillary driving force acting on that domain.
We note that the elastic modulus enters the poroelastic diffusion constant even when the gel is in its viscoelastic regime because the modulus controls the initial magnitude of the viscoelastic stress. In the viscoelastic regime, this stress does not persist indefinitely, but it governs the rate at which solvent redistributes while it is present. By contrast, in the elastic regime, the stress persists up to very long times as discussed below.
We note that most treatments of poroelasticity take the crosslinked gel to have an elastic (as opposed to viscoelastic) stress response [26, 12]. This indeed applies for times greater than the intrinsic viscoelastic crossover time of the gel, $\tau_{\mathrm{el}}$, which separates the early-time stress response, when polymer motions can relax an imposed stress, from the long-time elastic response [11, 22]. For $t\gg\tau_{\mathrm{el}}$, the shorter-time polymeric degrees of freedom have equilibrated, the crosslinked polymers behave elastically, and dissipation associated with stress redistribution is controlled primarily by solvent flow relative to a static polymer network. In this regime, the viscosity entering stress diffusion is well approximated by the Darcy viscosity, so that $\eta_{ps}\approx\eta_{D}$.
During coarsening, however, the gel is responding dynamically to stresses imposed by solvent being ejected from shrinking domains and accumulated in growing domains. These stresses act as localized perturbations to a gel that is otherwise close to its equilibrium swollen state. Because these perturbations are time dependent and spatially localized, they do not probe the static, long-time elastic response of the network, but instead drive transient polymer stress relaxation. Permanent crosslinking suppresses macroscopic flow but does not eliminate internal polymer dissipation; even dry covalently crosslinked networks exhibit viscoelastic stress relaxation at intermediate times due to strand-level and non-affine relaxation modes ($\sim$ a few hundred s), before reaching an elastic plateau [25]. For this reason, we envision the gel responding viscoelastically on the timescale of coarsening, before elastic arrest intervenes and halts the dynamics at long times (on the scale of hours). The viscoelastic response of the polymer, which applies even to permanently crosslinked gels, arises from the time-dependent, local strains in the gel due, for example, from the solvent being injected by a shrinking domain. These strains generate time-dependent polymer stresses that excite internal viscoelastic relaxation modes associated with frictional rearrangements of polymer segments between crosslinks. The local and transient polymer rearrangements contribute to stress relaxation, and the relevant dissipation expressed by the effective stress-relaxation viscosity $\eta_{ps}$ reflects both solvent–polymer friction and the intrinsic viscoelastic dynamics of the polymer-rich phase [22, 18].
Accordingly, we take the viscosity entering poroelastic diffusion during coarsening to be a poroelastic viscosity $\eta_{ps}$ that is distinct from the Darcy viscosity $\eta_{D}$. The Darcy viscosity characterizes solvent flow in response to pressure gradients in a static network, whereas $\eta_{ps}$ governs the coupled relaxation of pressure and polymer stress when the network itself is dynamically responding. This distinction must be made even in a permanently crosslinked gel. There too solvent transport generates strains on timescales comparable to polymer stress relaxation. In general, therefore, $\eta_{ps}\neq\eta_{D}$ for $t\ll\tau_{\mathrm{el}}$. We comment further on the implications of this distinction in the Results and Discussion.
When poroelastic diffusion governs the relaxation of pressure gradients, Eq. 1 can be written in the scaling form
$\dot{\lambda}\sim-\frac{k_{0}\sigma}{\eta_{D}\,\lambda\,\ell_{p}},$ (3)
where $\ell_{p}$ denotes the characteristic length over which pressure relaxes in the gel. The implications of this form, and the behavior of $\ell_{p}$ in different regimes, are discussed in the Results.
We consider two limiting regimes of poroelastic dissipation. (i) Fast poroelastic diffusion. If stress and pressure relax rapidly in the viscoelastic gel, the pressure gradient induced by capillary driving decays over a length scale of order the domain size. In this limit, coarsening follows kinetics analogous to the Lifshitz–Slyozov–Wagner case, with $\lambda\sim t^{1/3}$. (ii) Slow poroelastic diffusion. If poroelastic relaxation is slow, the pressure gradient decays over a poroelastic length that grows diffusively, $\ell_{p}\sim\sqrt{D_{p}\,t}$. Substituting this into Eq. 3 yields the scaling law
$\lambda\sim\left(\frac{k_{0}\,\sigma}{\eta_{D}\,D_{p}^{1/2}}\right)^{1/2}t^{1/4}.$ (4)
As above, shrinking and growing domains obey the same scaling by solvent conservation. In this slow-poroelastic regime, the coarsening law exhibits an explicit dependence on the gel modulus through the poroelastic diffusion constant, $D_{p}\sim G$, giving $\lambda\sim G^{-1/4}t^{1/4}$. The extent to which the prefactor $\sqrt{k_{0}\sigma/\eta_{D}}$ introduces additional modulus dependence is discussed further in the Results.
The use of the diffusion length $\ell_{p}=\sqrt{D_{p}\ t}$ for the slow poroelastic limit is only appropriate if $\ell_{p}<\lambda(t)$; relaxation of the gel on scales larger than the domain size $\lambda(t)$ is not relevant for solvent injection into the gel by a shrinking domain of size $\lambda$. Thus, the slow poroelastic limit crosses over to the fast limit at a crossover domain size. Equating $\ell_{p}\sim\lambda(t)$ yields a time-independent crossover length
$\lambda_{\times}\sim\,\frac{\sigma}{G}\,\frac{\eta_{ps}}{\eta_{D}},$ (5)
which magnifies the elastocapillary length $\sigma/G$ (which is of the order the mesh size of the gel) by a viscosity ratio. The crossover length can be large since the Darcy viscosity $\eta_{D}$ accounts for the solvent viscosity and friction with the instantaneous state of the polymers of the gel while the poroelastic viscosity in the viscoelastic regime of the gel, where coarsening takes place, accounts for polymer-polymer friction and local chain relaxation When the domain size $\lambda(t)<\lambda_{\times}$ the slow limit applies with coarsening scaling as $t^{1/4}$. In the opposite limit, the pressure gradients decay as $1/\lambda(t)$ which is appropriate in the fast poroelastic limit where the domain size scaling is $t^{1/3}$.
II.3 Elastic arrest of coarsening
So far, we have focused on the capillary driving force for coarsening. At early times, the coarsening gel is in the viscoelastic regime, where polymers respond locally and transiently to the stress imposed by solvent expelled from shrinking domains. In this regime, elastic stresses relax on times shorter than the coarsening time and there is no spatially long-ranged and temporally persistent elastic response; those transient stresses therefore act primarily as a source of dissipation, slowing—but not halting—the capillarity-driven coarsening dynamics.
At sufficiently long times, $t>\tau_{\mathrm{el}}$, the local and transient viscoelastic polymer relaxation modes coupled to the coarsening-induced stress have relaxed, and the polymer network responds collectively and elastically [11, 22]. Although the elastic modulus is a property of the network at all times, it becomes dynamically operative as a long-ranged elastic stress that resists solvent flow only in the elastic regime, when stresses generated by solvent transport can no longer relax locally and a temporally permanent and a spatially long-ranged stress profile is established in the gel. In an elastic gel, capillary forces are balanced by gel stress at the domain boundary, determined self-consistently by elastic deformation of the entire surrounding medium. Elastic arrest is therefore most transparently understood as a force balance at the boundary of the shrinking domain: the Laplace capillary pressure ($\delta p=\sigma/\lambda)$ is canceled by the elastic stress of the gel.
Crucially, the elastic stress at the domain boundary is not fixed a-priori. It must be determined by solving the elastic displacement problem in the gel surrounding the entire domain. For a spherical domain in an elastic solid [14], normal interfacial stress is of order
$\sigma_{\rm el}\;\sim\;G\,\frac{\delta\lambda}{\lambda}$ (6)
where $G$ is the shear modulus of the gel and $\delta\lambda$ is the elastic displacement of the interface. Physically, $\delta\lambda$ represents the instantaneous radial deformation of the gel at the domain boundary induced by the capillary pressure acting on the curved interface; in the elastic regime it is defined at all times as the solution of the elastic equilibrium equations with the instantaneous interfacial pressure $\delta p(t)\sim\sigma/\lambda(t)$ applied as a boundary condition, which in general is time dependent because $\lambda(t)$ evolves during coarsening.
Prior to arrest, the magnitude of $\delta\lambda$ is smaller than its long-time value because the long-ranged elastic stress field has not yet fully developed and therefore lags behind the capillary pressure. At long times, once the elastic stress has equilibrated spatially, the Laplace force balance at the domain boundary fixes $\delta\lambda$ by equating the elastic stress to the capillary pressure, $G(\delta\lambda/\lambda)\sim\gamma/\lambda$, which yields $\delta\lambda\sim\gamma/G$; thus, although the domain size $\lambda$ is mesoscopic, the interfacial displacement itself is set by a microscopic elastic length of order the gel mesh size. Because the elastic deformation extends over a region of size $\sim\lambda$, the domain size itself enters the solution: larger domains require elastic deformation over larger distances and therefore generate a different stress response. By contrast, the elastic stress at the solvent-rich domain boundary is not present in the viscoelastic, coarsening regime, where the unbalanced capillary pressure $\delta p$ is what drives solvent flow and coarsening.
We emphasize that elastic arrest is a kinetic effect that is relevant only for $t>\tau_{\mathrm{el}}$. Arrest therefore occurs when the coarsening time required to reach a domain size $\lambda$ becomes comparable to the intrinsic elastic crossover time:
$t(\lambda_{\rm arrest})\sim\tau_{\mathrm{el}}.$
Using the slow-poroelastic time-to-size relation of Eq. 4 ($\lambda^{4}\propto t$), we obtain the arrest size in the slow-diffusion regime,
$\lambda_{\rm arrest}^{(\rm slow)}\sim\Bigg(\,\frac{\eta_{ps}\,\eta_{el}}{\eta_{D}^{2}}\,\frac{k_{0}\,\sigma^{2}}{G^{2}}\,\Bigg)^{1/4},$ (7)
where we have expressed the elastic crossover time in terms of an effective elastic viscosity $\eta_{el}=G\tau_{\mathrm{el}}$, because viscosities scale as a modulus times a characteristic relaxation time. This form highlights that, in contrast to the coarsening kinetics, the arrest length depends only on ratios of viscosities and on elastic and interfacial material parameters.
In the fast-poroelastic regime, the corresponding arrest size is
$\lambda_{\rm arrest}^{(\rm fast)}\sim\Bigg(\,\frac{\eta_{el}}{\eta_{D}}\,\frac{k_{0}\,\sigma}{G}\,\Bigg)^{1/3}.$ (8)
III Results
In this section, we analyze the previous results to predict the scaling of the domain coarsening as a function of time and gel elastic modulus, as well as the modulus dependence of the arrest length. In addition, we show that this length can be magnified to a value far larger than the gel mesh size, by the viscosity ratios defined above. We present our analysis for two limiting cases of the dependence of the gel properties (permeability $k_{0}$ and interfacial tension with the solvent-rich phase $\sigma$) that enter the capillary-driven coarsening and eventual arrest. We keep the explicit dependence on the gel modulus $G$ since this is often changed and measured in experiments [9]. The first scenario we call the mesh-scaling gel where the polymer fraction in the gel is relatively small and the mesh size $\xi$ is the largest length in the homogeneous gel. The permeability in the Darcy equation has units of area and hence scales as $\xi^{2}$, while the tension (in units of $k_{B}T$) scales as $1/\xi^{2}$. The second scenario we call the melt-like gel where the polymer fraction in the gel is relatively large (order 1/2) and the mesh size is nanometric and can be of order the polymer persistence length and solvent molecule size. In contrast to the mesh-scaling gel with large pores which are the channels for solvent flow, the system may be conceptualized as somewhat swollen and crosslinked polymer melt. This is the limit that is applicable to the gels in [9] where the arrest length has been measured. For the melt-like gel, we keep $k_{0}$ and $\sigma$ as parameters which vary slowly with the gel mesh size, due to the involvement of the other length scales of the system.
We first present the scaling relations for the coarsening dynamics and the arrest length as functions of the three viscosities defined above: The Darcy viscosity $\eta_{D}$, the poroelastic viscosity for stress diffusion in the viscoelastic regime of the polymers $\eta_{ps}$, and the viscosity $\eta_{el}=G\ \tau_{el}$ related to the time scale $\tau_{el}$, at which the viscoelastic regime crosses over to the elastic regime. Afterwards, we rewrite the scaling relations as functions of time scales associated with the Darcy and poroelastic viscosities, guided by the fact that a viscosity is the product of a modulus and a time scale. This effective modulus characterizes the short-time resistance of the polymer network to deformation by solvent flow. For scaling purposes, the distinction between this viscoelastic modulus and the long-time elastic modulus $G$ affects only prefactors and is therefore neglected since both are governed by the same crosslinked, network structure [23, 11]. We therefore write $\eta_{D}=G\ \tau_{D}$ and $\eta_{ps}=G\ \tau_{ps}$ where the dependence on $G$ reflects that the solvent flow, is impeded by friction with the gel even in its viscoelastic regime. The results for the scaling relations are shown in Tables 1 and 2.
Table 1: Coarsening laws and kinetic-arrest lengths with dimensional prefactors in the fast and slow poroelastic regimes, for the two gel models with independent viscosities.
Gel Model Regime Coarsening law $\lambda(t)$ Arrest length $\lambda_{\rm arrest}$
(i) Mesh-scaling gel: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim\left(\frac{k_{B}T}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/12}\left(\frac{\eta_{ps}}{\eta_{D}^{2}}\right)^{1/4}t^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
(ii) Melt-like gel: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G}\,\frac{\eta_{ps}}{\eta_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
Table 2: Same as Table 1, but rewriting all viscosities as products of an elastic modulus and characteristic time scales.
Gel Model Regime Coarsening law $\lambda(t)$ Arrest length $\lambda_{\rm arrest}$
(i) Mesh-scaling gel + $\eta\sim G\tau$: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
(ii) Melt-like gel + $\eta\sim G\tau$: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{G}\,\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
We note that in Table 2 where we have extracted the maximal dependence on the elastic modulus $G$ that is physically reasonable, the fast and slow coarsening for the mesh-scaling gel both scale as $1/{G}^{1/3}$ with respective time dependence of $t^{1/3}$ and $t^{1/4}$ respectively. For the melt-like gel, appropriate for high polymer volume fractions where the porosity and interfacial tension depend only weakly on $G$, the fast coarsening law exhibits the same scaling as for the mesh-scaling gel. On the other hand, the slow coarsening law still scales as $t^{1/4}$ but with a coefficient that now scales as $1/\sqrt{G}$. This is also the modulus scaling of the arrest length in the slow poroelastic limit.
IV Discussion
Interpretation of mesh-scaling gel: scaling of domain coarsening with time and elastic modulus: The only length scale in the problem is the gel mesh size $\xi\sim 1/G^{1/3}$. The fast poroelastic diffusion predicts coarsening which scales as $t^{1/3}$ while slow poroelastic diffusion coarsening scales as $t^{1/4}$. If we further take the viscosities to all scale as the modulus $G$ and a different effective relaxation times as in Table 2, the dependence on the elastic modulus in all cases, for both the coarsening law and kinetic arrest lengths scales as the mesh size, proportional to $1/G^{1/3}$. The coarsening lengths which are power laws in the time, naturally contain a characteristic time scale. For the fast coarsening this is the relaxation time of the Darcy flow $\tau_{D}$ appropriate for solvent flow against the friction of the polymers in their static configurations. The slow coarsening which proceeds via poroelastic stress diffusion in the gel with the flowing solvent, has a characteristic time proportional to the ratio $\tau_{D}^{2}/\tau_{ps}$. Since $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains induced by the solvent flow, in systems with a high polymer fraction, we expect $\tau_{ps}\gg\tau_{D}$ so the characteristic time scale of the slow coarsening is small compared to $\tau_{D}$ itself. The fact $\tau_{ps}\gg\tau_{D}$ is the reason that the crossover length scale $\lambda_{\times}$ in Eq. 5 is magnified from the mesh scale to the mesoscale. Nevertheless, the “fast” and “slow” nomenclature refers to the coarsening exponents rather than the characteristic time scales: even when the slow poroelastic mode is governed by a shorter characteristic time scale, it still yields slower growth in time ($t^{1/4}$ versus $t^{1/3}$).
Melt-scaling gel: scaling of domain coarsening with time and elastic modulus: For this system, we keep the porosity and interfacial tension as parameters that vary only slowly with the modulus $G$, since for large polymer fraction gels (the melt-scaling limit), additional length scales related to the Kuhn length and solvent size can determine these parameters. However, despite this, the crosslinked, melt-scaling gel is still well characterized by its elastic modulus $G$ which remains the dominant parameter controlling both stress transmission and arrest. In this regime, fast poroelastic diffusion predicts coarsening which scales as $t^{1/3}$ while slow poroelastic diffusion coarsening scales as $t^{1/4}$. If we further take the viscosities to all scale as the product of the gel modulus $G$ and different effective relaxation times as in Table 2, for fast poroelastic coarsening, the dependence on the elastic modulus in all cases, for both the coarsening and the arrest lengths scale with $1/G^{1/3}$.
On the other hand, for slow poroelastic coarsening of the melt-like gel, we stress that both the coarsening and arrest lengths of Table 2 scale as $1/\sqrt{G}$. The arrest length is magnified from its microscopic or mesh-size value by the ratio of the time scales $\tau_{el}\gg\tau_{D}$ for fast coarsening and $\tau_{el}\tau_{ps}\gg\tau_{D}^{2}$ for slow coarsening. These large ratios reflect the separation between solvent transport and polymer stress relaxation processes: $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains activated by the solvent flow while $\tau_{D}$ is related to solvent flow in the presence of polymer friction in an essentially static configuration. The longest time scale is $\tau_{el}$ which is the crossover from viscoelasticity (activated by solvent flow) to elastic response when further coarsening is kinetically arrested by elastic stress acting over a region of order $\lambda$. We emphasize that the arrest discussed here refers to the breakdown of capillarity-driven, pressure-mediated coarsening; while other, much slower mechanisms of structural evolution cannot be excluded in principle, they lie outside the scope of the present theory.
Comparison with experiment: The melt-like gel predictions may be applicable to the experiments of [9] where the gels are covalently crosslinked and the polymer fraction is relatively large (of order 1/2). Unfortunately, the coarsening kinetics has not yet been measured in these system. Our predictions for slow poroelastic diffusion for the melt-like gel, for scaling of the coarsening length with $t^{1/4}$ and with elastic modulus as $1/\sqrt{G}$, do agree with the numerical model calculations in [9] but the significance of this agreement is not clear since their model is based on Cahn-Hilliard dynamics with a constant mobility while our scaling predictions originate in the Darcy driven solvent flow. Although a $t^{1/4}$ growth law is also known [3] to arise in coarsening via surface–diffusion–limited coarsening where material transport occurs along connected interfaces, this mechanism is not applicable to the present system since observability of coarsening in the experiments (see the SI of [9] ) depends on the elastic modulus of the bulk material. In addition, in the three softest gels of [9] the solvent-rich domains are distinct and unconnected.
The primary prediction of our scaling model that does agree with the measurements in [9] is that the arrest length-scale varies for the melt-like gel for slow poroelastic diffusion, as $1/G^{1/2}$. This agreement indicates that the gels in these polymer-rich systems are most probably not in the mesh-scaling or the fast poroelastic diffusion regimes, where the arrest length scale is predicted to vary as $1/G^{1/3}$. Our scaling relations also predict the dependence of the arrest-length on the various viscosities and future measurements of this behavior (see below) would be of interest.
We now address the very slow drainage of solvent at time scales of about a day. For that process, also driven by capillarity involving the solvent in the gel and the excess solvent reservoir, the solvent flow is slow and we expect that the macroscopic drainage to depend on poroelastic diffusion, but with Darcy flow related to the solvent viscosity $\eta_{s}$ and not $\eta_{ps}$ since the polymers have relaxed all stresses on the macroscopic time scale. We thus estimate that the macroscopic time for drainage scales as $R^{2}/D_{D}$ where $R$ is the macroscopic sample size and the Darcy poroelastic diffusion constant $D_{D}$ scales as $k_{0}G/\eta_{s}$. For an order of magnitude numerical estimate, we take the permeability $k_{0}\sim\xi^{2}\sim(k_{B}\ T/G)^{2/3}$ and $\eta_{S}$ to be of the order of magnitude of the water viscosity, $10^{-3}\text{Pa~s}$. For $G\sim 100\text{kPa}$ as a typical rigidity [9], we find for a 1cm sample, a drainage time of the order of one day.
We conclude by emphasizing that the scope of our scaling picture is more general than its application to specific gels of Ref. [9], which we interpret as representative of melt-like gels. Our predictions for the mesh-scaling gels can be tested by working at significantly lower polymer volume fractions, where the gel enters the mesh-scaling regime and distinct coarsening laws and arrest scales are expected.
Even more generally, our framework applies to both permanently and non-permanently crosslinked gels. In non-permanently crosslinked gels, where the homogeneous swollen state already exhibits viscoelastic stress relaxation, the effective viscosities introduced here can be directly related to the linear rheological properties of the uniform gel. However for permanently crosslinked gels, the viscosities appearing in our theory are intrinsically non-equilibrium quantities, controlled by polymer chain relaxation in the presence of solvent flow, such as that induced by shrinking solvent-rich domains. In such systems, these effective viscosities may be independently probed through measurements of the kinetics of solvent uptake or release of a macroscopic gel in contact with a solvent reservoir. This allows experimental tests of the dynamical ingredients of our theory across a wide range of gel architectures. Such measurements would allow a direct comparison between macroscopic poroelastic response and microscopic coarsening dynamics.
V Acknowledgements
We are grateful to Eric Dufresne and Carla Fernandez-Rico for discussions about their experiments and to David Andelman, Haim Diamant, Amit Kumar, Biplab Mandal, Manu Mannattil, and David Zwicker for insights into their theoretical models.
Appendix A Onuki-Puri numerical calculations
A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction. In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus. They found that for an uncrosslinked gel (their effective crosslink density, is equivalent to the modulus $G$ in units of $k_{B}T$) coarsening was governed by a 1/3 power law. When varying degrees of crosslinking were included, the coarsening was governed by an effective power law where the exponent depended on the crosslink density (see their Fig. 2 and caption). They also observed the onset of elastic arrest with a time that depended on the value of the modulus. The dynamics of their model only included polymer-solvent friction and the gel had no viscoelastic regime and could not include Darcy transport, poroelastic stress diffusion, and viscoelastic stress buildup, all important elements of our approach and its relevance to the experiments in [9]. In their simulations, the effective network–solvent friction includes an explicit Stokes prefactor ($6\pi$), which enters directly into the arrest criterion. As a result, the arrested length can be enhanced by roughly an order of magnitude relative to the naive elastocapillary estimate $\sigma/G$, even in the absence of an additional large control parameter. By contrast, in our scaling theory the viscosities enter only through ratios (see the Tables below), so such numerical prefactors cancel. The remaining microscopic input enters the arrest length only to a fractional power, rendering the predicted mesoscopic magnification robust and weakly sensitive to microscopic details.
Appendix B Curk and Luijten model
A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix. Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects, and does not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach (see Appendix for further discussion). They predict an initial regime of growth followed by elastic arrest at short times, an intermediate viscoelastic coarsening regime with linear-in-time growth, and a late-time crossover back to LSW coarsening with domain size scaling as $t^{1/3}$. The late-time return to LSW scaling is not observed in PDMS gels, where the long-time (hours) state is one of elastic arrest, indicating a distinct rate-limiting physical mechanism.
Appendix C Unified diffusion-based description of coarsening in fluids and gels
In this Appendix we present a unified derivation of the coarsening dynamics for (i) phase separation in a binary fluid mixture of small molecules, and (ii) solvent-rich domains embedded in a polymer gel. Although the microscopic transport mechanisms differ in the two cases, both can be formulated in terms of a diffusion equation for the chemical potential or equivalently, for the pressure; both use the capillary (Laplace pressure jump) boundary condition at the domain interface. This unified formulation clarifies which approximations underlie the standard LSW result for simple fluids and how these are modified in gels by slow poroelastic transport.
C.1 Interfacial flux and capillary boundary condition
We consider a spherical solvent-rich domain of radius $\lambda(t)$ that is shrinking in time. Let $N$ denote the number of transported molecules (B molecules in a binary fluid, or solvent molecules in a gel) inside the domain. Mass conservation at the moving interface gives
$\frac{dN}{dt}=-JS,$ (9)
where $J$ is the outward flux of transported molecules evaluated at the domain boundary and $S=4\pi\lambda^{2}$ is the surface area of the domain. The flux is normal to the interface and is controlled by the local driving force acting at the boundary.
To relate $\dot{N}$ to the interface velocity $\dot{\lambda}$, we note that a change in radius by $d\lambda$ changes the domain volume by $dV=S\,d\lambda$. The net number of transported molecules lost from the domain when a volume $dV$ is removed is $\Delta c\,dV$, where
$\Delta c\equiv c_{\rm in}-c_{\rm out}$ (10)
is the concentration difference between the inside of the domain and the surrounding phase. Therefore,
$\dot{N}\sim\Delta c\,S\,\dot{\lambda}.$ (11)
[…]
$\Delta c\,\dot{\lambda}\sim-J.$ (12)
The flux can be written in terms of the chemical potential $\mu$ of the transported species as
$J=-M\nabla\mu,$ (13)
where $M$ is a mobility.
C.2 Mobility in binary fluids and gels
The mobility $M$ appearing in the interfacial flux relation, $J=-M\nabla\mu$, originates in the microscopic mechanism by which the transported species moves in response to a chemical potential gradient. Its physical meaning differs in the two systems considered here.
Binary fluid (small molecules).
For a mixture of small molecules, transport occurs by ordinary molecular diffusion. The mobility is related to the mutual diffusion constant $D$ via the Einstein relation,
$M\sim D\,c_{B},$ (14)
where $c_{B}$ is the number density of the transported species. Equivalently, using $\nabla\mu=\nabla p/c_{B}$, the flux may be written as $J\sim-D\nabla c_{B}$. In this case, the mobility is due to the rapid diffusive motion of individual molecules in the surrounding fluid.
Gel (poroelastic transport).
In a gel, transport of solvent is instead controlled by permeation through a deformable polymer network. The appropriate constitutive relation is Darcy’s law,
$\vec{v}=-\frac{k_{0}}{\eta_{D}}\nabla p,$ (15)
where $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity associated with solvent flow impeded by friction with the polymer network. The solvent flux is $J=c_{s}v$, where $c_{s}$ is the solvent concentration, so that the mobility entering the flux relation is
$M\sim\frac{k_{0}\,c_{s}}{\eta_{D}}.$ (16)
In contrast to the binary-fluid case, the mobility here is controlled by hydrodynamic permeation of solvent through the network rather than by molecular diffusion.
Thus, while the diffusion equation for the pressure field has the same mathematical form in both systems, the mobility MM—and therefore the physical origin of dissipation—is fundamentally different in binary fluids and in gels.
The capillary driving force originates in the curvature dependence of the chemical potential at the interface. For a spherical domain of radius $\lambda$, the chemical potential at the interface exceeds that of a flat interface by
$\delta\mu=\frac{\sigma}{c\,\lambda},$ (17)
where $\sigma$ is the interfacial tension and $c$ is the concentration of the transported species in the domain.
C.3 Relation of mechanical and osmotic pressures
In an incompressible soft-matter system, local variations in solvent concentration generate osmotic stresses that tend to drive solvent flow. Because local volume changes are forbidden, these stresses give rise to spatial variations of a mechanical pressure field, which enforces incompressibility and appears in the force balance governing solvent motion. This mechanical pressure is a physical stress that adjusts so as to balance all forces acting on the material under the incompressibility constraint.
Gradients of chemical potential are directly related to gradients of osmotic pressure, since the osmotic pressure measures the change in free energy associated with locally adding or removing material. Elastic stresses arising from deformation of the polymer network also contribute to the mechanical pressure, but they do so through the network’s mechanical response to imposed stresses rather than through a direct coupling to composition. As a result, elastic stresses influence the pressure field indirectly, by determining how osmotic stresses are transmitted through the network.
In regimes where the polymer network can relax locally — corresponding to the viscoelastic or flowing regime relevant on the coarsening time scale— elastic stresses adjust rapidly and do not build up over large distances. As a result, the pressure gradients that drive solvent flow are set primarily by gradients of osmotic pressure.
By contrast, at sufficiently long times the network responds elastically over length scales comparable to the domain size, and elastic stresses accumulate over these large scales. In this elastic regime, further solvent transport requires deformation of the entire domain, and the resulting pressure gradients are determined by force balance across the domain rather than by local osmotic gradients alone. This buildup of elastic stress leads to elastic arrest of coarsening.
Thus, while elastic stresses are essential for determining the arrest of coarsening at long times, on the coarsening time scale itself the polymer network behaves viscoelastically, and elastic stresses relax locally and do not contribute significantly to the pressure gradients that drive flow.
Capillary stresses enter separately as boundary conditions at interfaces: the Laplace law fixes the pressure jump across the interface between the solvent-rich domain and the gel and thereby sets the magnitude of the pressure difference that must be accommodated by the pressure field in the surrounding material.
In summary, while the mechanical pressure field includes osmotic, elastic, and capillary contributions, gradients of chemical potential may be equivalently expressed in terms of pressure gradients because osmotic stresses provide the fundamental thermodynamic driving force, with transient and local elastic stresses controlling how this driving force is transmitted in the viscoelastic regime of the gel.
C.4 Relation between chemical potential and pressure gradients.
For an incompressible binary system, we consider a free-energy density $f(c)$ that depends on the local concentration $c$ of the transported species (the solvent in a gel, or component $B$ in a small-molecule mixture). The chemical potential is
$\mu=\frac{df}{dc_{B}}.$ (18)
For a system containing $N$ particles in a volume $V$ with concentration $c=N/V$ and free-energy density $f(c)$, the total free energy is
F=Vf(c).F=V\,f(c).
The osmotic pressure is defined as the force per unit area opposing a change in volume at fixed particle number,
$\Pi\equiv-\left(\frac{\partial F}{\partial V}\right)_{N}.$
Using dc=(c/V)dVdc=-(c/V)\,dV, one finds
$\Pi=c\frac{df}{dc}-f(c).$
Identifying the chemical potential as $\mu=df/dc$, this can be written as
$\Pi=\mu\,c_{B}-f(c_{B}).$
To make this relation explicit, we differentiate
$\Pi=\mu\,c-f(c)$
with respect to the state variables at fixed temperature. Using the chain rule, this gives
$d\Pi=\mu\,dc+c\,d\mu-df.$
By definition of the chemical potential,
$\mu\equiv\frac{df}{dc},$
so that
$df=\mu\,dc.$
Substituting this expression into the equation above, the terms proportional to $dc$ cancel, leaving
$d\Pi=c\,d\mu.$
This shows that gradients in chemical potential are equivalently expressed as gradients in osmotic pressure, which provide the mechanical driving force for transport. Therefore, at fixed temperature,
$\nabla\Pi=c\,\nabla\mu.$ (19)
We recall that $c$ is the concentration of the transported species (the B molecules in the AB small-molecule mixture or the solvent molecules in the solvent-gel system) outside the coarsening solvent-rich (or B molecule rich) domain.
In an incompressible system, the mechanical pressure (outside the solvent-rich domain) that drives transport differs from the osmotic pressure only by a reference constant (and by elastic or capillary contributions treated separately through force balance and (Laplace) boundary conditions at the solvent-rich domain interface). Consequently, pressure gradients may be identified with osmotic pressure gradients, $\nabla p=\nabla\Pi$, which gives the thermodynamic relation
$\nabla\mu=\frac{1}{c}\,\nabla p.$ (20)
This allows the flux to be written equivalently as
$J=-\frac{M}{c}\,\nabla p.$ (21)
The capillary boundary condition fixes the mechanical pressure at the domain surface through the Laplace (osmotic) pressure jump $\delta p$,
$p(r=\lambda,t)=\delta p\equiv\frac{\sigma}{\lambda}.$ (22)
C.5 Diffusion equation for the pressure field
We show now that the mechanical pressure field satisfies a diffusion equation,
$\frac{\partial p}{\partial t}=D\nabla^{2}p,\qquad r\geq\lambda,$ (23)
This result follows from the combination of (i) the expression for the solvent flux, written in terms of pressure gradients, and (ii) local conservation of solvent.
The outward number flux $J$ of the transported species is proportional to the driving-force gradient,
$J=-M\nabla\mu,$ (24)
where $M$ is the mobility. Using the incompressible thermodynamic relation $\nabla\mu=\nabla p/c_{B}$, this may be written as
$J=-\frac{M}{c_{B}}\,\nabla p.$ (25)
Let $c(\mathbf{r},t)$ be the number density of the transported species in the medium outside the domain (the A-rich phase for a binary fluid, or the polymer-rich phase for a gel). Local conservation implies
$\frac{\partial c}{\partial t}+\nabla\!\cdot J=0.$ (26)
[…]
$\frac{\partial c}{\partial t}=\nabla\!\cdot\!\big(M\nabla\mu\big).$ (27)
For small departures from equilibrium in the surrounding medium outside the domain, where the pressure and composition fields adjust diffusively to the interfacial boundary condition, variations in chemical potential are proportional to composition variations,
$\delta\mu\simeq\left(\frac{\partial\mu}{\partial c}\right)\delta c,$ (28)
[…]
$\frac{\partial c}{\partial t}\simeq D\,\nabla^{2}c,\qquad D\equiv M\left(\frac{\partial\mu}{\partial c}\right),$ (29)
where $D$ is the appropriate diffusivity (molecular or poroelastic, depending on the system).
[…]
$\frac{\partial p}{\partial t}=D\nabla^{2}p,\qquad r\geq\lambda,$ (30)
with boundary conditions
$p(r=\lambda,t)=\delta p,\qquad p(r\to\infty,t)=0,$ (31)
and initial condition $p(r,t=0)=0$. Thus the diffusive spreading of $p$ (and the associated length scale $\ell\sim\sqrt{Dt}$) follows directly from a gradient-driven flux and local conservation.
In a binary fluid mixture of small molecules, $M\sim D\,c_{B}$ and the diffusion equation reduces to ordinary molecular diffusion of composition (or chemical potential) in the continuous phase. In a gel, the same mathematical structure arises from poroelastic transport: Darcy’s law together with force balance and network elasticity leads to a diffusion equation for pressure with $D=D_{p}\sim k_{0}G/\eta_{ps}$ (see[26])
The physical meaning of the diffusion constant $D$ differs in the two cases considered here:
  • Binary fluid (small molecules): $D$ is the molecular mutual diffusion constant of the B molecules in the A-rich phase.
  • Gel (poroelastic transport): $D=D_{p}$ is the poroelastic diffusion constant,
    $D_{p}\sim\frac{k_{0}G}{\eta_{ps}},$ (32)
    where $k_{0}$ is the permeability of the gel, $G$ its elastic modulus, and $\eta_{ps}$ the viscosity associated with the viscoelastic response of the polymer–solvent system.
C.6 Exact solution around a spherical domain
[…]
$p(r,t)=\delta p\ \frac{\lambda}{r}\mathrm{erfc}\!\left(\frac{r-\lambda}{2\sqrt{Dt}}\right),\qquad r\geq\lambda.$ (33)
The radial pressure gradient evaluated at the domain boundary is
$\left.\frac{\partial p}{\partial r}\right|_{r=\lambda}=-\delta p\left(\frac{1}{\lambda}+\frac{1}{\sqrt{\pi Dt}}\right).$ (34)
This expression shows explicitly how the interfacial pressure gradient evolves in time as the pressure field spreads into the surrounding medium.
C.7 Binary fluid: fast molecular diffusion (LSW limit)
[…]
$\left.\frac{\partial p}{\partial r}\right|_{r=\lambda}\simeq-\frac{\delta p}{\lambda}.$ (35)
[…]
C.8 Gel: poroelastic diffusion and crossover regimes
[…]
$\ell_{p}(t)=\sqrt{D_{p}t},$ (36)
the interfacial pressure gradient can be written as
$\left.\nabla p\right|_{r=\lambda}=-\delta p\left(\frac{1}{\lambda}+\frac{1}{\sqrt{\pi}\,\ell_{p}}\right).$ (37)
This expression exhibits two limiting regimes:
  1. 1.
    Slow poroelastic diffusion ($\ell_{p}\ll\lambda$):
    $\left.\nabla p\right|_{r=\lambda}\sim-\frac{\delta p}{\ell_{p}},$
    corresponding to a transient buildup of the interfacial pressure gradient.
  2. 2.
    Fast poroelastic diffusion ($\ell_{p}\gg\lambda$):
    $\left.\nabla p\right|_{r=\lambda}\sim-\frac{\delta p}{\lambda},$
    where the interfacial gradient has saturated and no longer depends on time.
[…]
$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c_{B}}\,\left.\nabla p\right|_{r=\lambda},$ (38)
so that the coarsening kinetics are entirely controlled by the pressure gradient evaluated at the domain boundary.
C.9 Coarsening laws
Inserting the mobility $M$ into the interfacial evolution law $\Delta c\,\dot{\lambda}\sim-M\,\left.\nabla\mu\right|_{r=\lambda}$ (or equivalently $\Delta c\,\dot{\lambda}\sim-(M/c_{B})\,\left.\nabla p\right|_{r=\lambda}$) immediately yields the coarsening kinetics, because the dynamics are controlled by the local interfacial gradient evaluated at $r=\lambda$.
C.9.1 Binary fluid (small molecules; fast molecular diffusion / LSW).
For a binary fluid of small molecules, transport in the continuous phase is due to ordinary molecular diffusion with diffusion constant $D$. As shown above, the pressure field outside a shrinking domain satisfies a diffusion equation and, because molecular diffusion is fast compared to the coarsening dynamics, the pressure profile rapidly reaches its long-time limit. Equivalently, the pressure satisfies Laplace’s equation outside the domain, with the capillary boundary condition $p(r=\lambda)=\delta p\sim\sigma/\lambda$.
In this limit, the interfacial pressure gradient is time independent and scales as
$\left.\nabla p\right|_{r=\lambda}\sim\frac{\delta p}{\lambda}\sim\frac{\sigma}{\lambda^{2}}.$
Using the interfacial evolution law $\Delta c\,\dot{\lambda}\sim-(M/c)\left.\nabla p\right|_{r=\lambda}$ with $M\sim Dc$ for molecular diffusion then gives
$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c}\left.\nabla p\right|_{r=\lambda}\sim-\frac{1}{\Delta c}\,D\,\frac{\sigma}{\lambda^{2}}.$ (39)
Integrating this equation yields the standard LSW coarsening law $\lambda(t)\sim(D\sigma\,t/\Delta c)^{1/3}$.
C.9.2 Gel (poroelastic; slow poroelastic diffusion, $\ell_{p}\ll\lambda$).
The interfacial pressure is set by capillarity, $\delta p\sim\sigma/\lambda$, and Darcy flow gives $M\sim k_{0}c_{s}/\eta_{D}$ with $\nabla\mu=(1/c_{s})\nabla p$. Heree, $c_{s}$ is the solvent concentration in the gel where it is transported by coarsening of the solvent-rich domain. For slow poroelastic spreading, the exact diffusion solution yields $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}\sim(\sigma/\lambda)/\ell_{p}$, so that
$\dot{\lambda}\sim-\frac{1}{\Delta c_{s}}\,\frac{k_{0}c_{s}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\frac{c_{s}}{\Delta c_{s}}\,\frac{\sigma}{\lambda\,\ell_{p}(t)},\qquad\ell_{p}(t)=\sqrt{D_{p}t},\ \ D_{p}\sim\frac{k_{0}G}{\eta_{ps}}.$ (40)
For scaling purposes, the dimensionless ratio $c_{s}/\Delta c_{s}$ is treated as a constant of order unity and we omit its numerical value here.
C.9.3 Gel (poroelastic; fast poroelastic diffusion, $\ell_{p}\gg\lambda$).
Once poroelastic diffusion has spread over distances comparable to the domain size, the interfacial gradient saturates to $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda\sim\sigma/\lambda^{2}$, and the evolution law becomes
$\dot{\lambda}\sim-\,\frac{k_{0}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\,\frac{\sigma}{\lambda^{2}}.$ (41)
This has the same $\lambda^{-2}$ structure as LSW, but with the molecular mobility $D$ replaced by the Darcy permeation mobility $k_{0}c_{s}/\eta_{D}$, and with the additional dimensionless factor $c_{s}/\Delta c_{s}$ which we omitted for scaling purposes.
Depending on whether the interfacial pressure gradient has reached its saturated value $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda$ or is still building up in time as $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}(t)$, one recovers either the standard LSW coarsening law or the poroelastic coarsening laws discussed in the main text.

### v2

@@
 Abstract
-Recent experiments on crosslinked gels quenched from solvent-rich to solvent-poor conditions, show solvent-rich domains surrounded by gel-rich regions that coarsen followed by kinetic arrest at micron scales that persists for hours before full drainage to macroscopic equilibrium occurs on day timescales.
-Motivated by this, we present a general model for both coarsening and eventual arrest of pressure-driven coarsening in gels with solvent flow driven by interfacial tension of the solvent-rich domains and the gel.
-In gels in their viscoelastic time regime, this capillary force is dissipated by solvent transport through crosslinked polymers where the polymers develop transient elastic stresses due to solvent flow.
-At longer times, the long-ranged, elastic response of the gel provides a deterministic force which balances the capillary force thus arresting pressure-driven coarsening at a stiffness-dependent length.
-For a melt-like, polymer-rich gel, the coarsening length $\sim t^{1/4}$ where $t$ is the time, with an amplitude $\sim G^{-1/2}$, where $G$ is the shear modulus.
-The arrest length in this regime also scales $\sim G^{-1/2}$.
-For low polymer fractions where the dominant length scale is the mesh size, the coarsening length can scale as $t^{1/3}$ varying with $G^{-1/3}$ (as does the arrest length).
-Our predictions for the arrest-length scaling with $\sim G^{-1/2}$ for the melt-like gel are consistent with the measurements.
+Recent experiments on crosslinked gels quenched from solvent-rich to solvent-poor conditions show solvent-rich domains embedded in a gel-rich matrix.
+These domains coarsen and then undergo kinetic arrest at micron scales for hours, before macroscopic drainage to equilibrium over an even longer time.
+Motivated by these observations, we develop a minimal model that couples capillarity-driven Darcy permeation to the viscoelastic-to-elastic crossover of the polymer network.
+In the viscoelastic regime, the Young–Laplace boundary traction (due to surface tension) at curved solvent–gel interfaces generates an isotropic pressure gradient in the solvent pores of the gel that drives solvent flow and coarsening.
+In the elastic regime, the same interfacial traction is balanced by a long-ranged deviatoric (non-isotropic) stress field, and the elastic equilibrium solution yields a spatially uniform pore pressure outside a domain, eliminating the Darcy driving force and arresting coarsening.
+Using the kinetic criterion $t(\lambda_{\rm arrest})\sim\tau_{\rm el}$, we predict stiffness-dependent growth laws and arrest lengths.
+For melt-like, polymer-rich gels, $\lambda(t)\sim G^{-1/2}t^{1/4}$ and $\lambda_{\rm arrest}\sim G^{-1/2}$.
+For low polymer fractions where the mesh size controls transport, $\lambda(t)\sim G^{-1/3}t^{1/3}$ and $\lambda_{\rm arrest}\sim G^{-1/3}$.
+The predicted $G^{-1/2}$ arrest scaling for melt-like gels is consistent with measurements.
 I Introduction
@@
 The temperature is then lowered, and the equilibrium solvent fraction within the gel decreases.
-Once the gel comes to equilibrium—which only occurs macroscopically after a day or two—the gel and the solvent exhibit phase separation.
+Once the gel comes to true equilibrium — which only occurs after many hours — the gel and the solvent exhibit macroscopic phase separation.
 At its lower temperature state, the gel initially contains excess solvent due to its initial high-temperature preparation.
@@
 Strikingly, in these experiments the coarsening is observed to undergo long-lived kinetic arrest at submicron or micron length scales, depending on the rigidity of the cross-linked gel.
-This mesoscale structure remains stable for several hours until, on the scale of days, all excess solvent eventually drains from the solvent-rich domains into a reservoir.
+This mesoscale structure remains stable for several hours until, after an even longer time, all excess solvent eventually drains from the solvent-rich domains into a reservoir.
 The gel then becomes homogeneous at its low-temperature state with a lower solvent fraction than its prepared state.
@@
 However, the paper states that for the more rigid gels, no coarsening was observed on the time scale of the measurements, while for the three softest gels, coarsening of the domains was observed.
-A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix.
+A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\gamma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix (see Appendix A.
 Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects.
@@
 In their numerical solutions the onset of elastic arrest occurred at a time that depended on the value of the modulus (see Appendix for further discussion).
-A more recent theory of phase separation in elastic networks [20], based on coupled chemical-potential gradients and elastic stress, concluded that local elasticity alone cannot arrest coarsening, and that non-local elastic interactions are required.
-By contrast, our approach focuses explicitly on solvent flow through the gel coupled to its poroelastic response.
-By treating solvent permeation (Darcy flow) and elastic back-reaction directly—rather than incorporating elasticity solely through constitutive stress–energy relations—we demonstrate that elastic arrest of coarsening can occur even for linearly elastic gels, as explained below.
+A more recent theory of phase separation in elastic networks [20], based on coupled chemical-potential gradients and elastic stress, concluded that local elasticity alone cannot arrest coarsening, and that non-local elastic interactions are required (see Appendix B.
+By contrast, our approach focuses explicitly on solvent flow through the gel coupled to the polymers’ poroelastic response.
+By treating solvent permeation in the gel (Darcy flow) and the elastic back-reaction directly—rather than incorporating elasticity solely through constitutive stress–energy relations—we demonstrate that elastic arrest of coarsening can occur even for linearly elastic gels, as explained below.
 Other theories [21, 16, 10] proposed to explain these phenomena using equilibrium models, unrelated to coarsening and poroelastic flow.
@@
 Here, we present a unified scaling theory of both domain coarsening and elastic arrest in crosslinked gels which exhibit crosslink-controlled elastic behavior at long times and a viscoelastic response activated by solvent flow through the network at intermediate times.
-Darcy flow [26, 2, 13] of the solvent in the gel occurs at time scales for which poroelastic dissipation, acting within the viscoelastic regime of the polymer network, controls the rate of coarsening, while the onset of long-time elastic response governs kinetic arrest [8, 17, 26].
-When the characteristic coarsening time becomes comparable to the relaxation time separating flowing, viscoelastic regime from the solid-like elastic regime, further capillary-driven domain evolution is arrested by the development of long-range elastic stress in the gel.
-At the level of the Laplace force balance at the solvent-rich domain boundary, this elastic stress cancels the capillary driving force and halts further solvent flow.
+Solvent flow (through “Darcy’s law”) is driven by the pressure gradients within the solvent pores of the gel [25, 2, 13].
+It occurs at time scales for which poroelastic dissipation, acting within the viscoelastic regime of the polymer network, controls the rate of coarsening.
+In contrast, the onset of long-time elastic response governs kinetic arrest [8, 17, 25], since as discussed below, the elastic equilibrium is that of a shear stress with no pressure gradient in the gel and hence no solvent flow or coarsening.
+When the characteristic coarsening time becomes comparable to the relaxation time separating the flowing, viscoelastic regime from the solid-like elastic regime of the crosslinked polymers, further capillary-driven domain evolution is arrested by the development of long-range elastic stress in the gel.
+At the level of the Laplace force balance at the solvent-rich domain boundary, the shear stress balances the capillary driving force; there is no pressure gradient in the solvent pores of the gel and no driving force for further solvent flow.
 The resulting arrested domain size is therefore not set directly by microscopic gel length scales, but by a kinetic criterion and is amplified by a magnification factor arising from the large separation between the poroelastic time governing solvent transport and the much longer time required for the onset of elastic response.
@@
 Thus, instead of equalities, we use the convention $\sim$ to mean “scales as”.
-The theory predicts a coarsening exponent of $1/4$ for slow, poroelastic diffusion and the arrest size of the domains scales as $1/\sqrt{G}$; the latter agrees with the experiments.
+The theory predicts that the coarsening length scales as $t^{1/4}$ for slow, poroelastic diffusion and that the arrest size of the domains scales as $1/\sqrt{G}$; the latter agrees with the experiments.
 Finally, we note here that while the explanations below consider the case of shrinking, solvent-rich domains, our model and its scaling predictions apply to both shrinking and growing domains, which are of course related by conservation of solvent.
@@
 At early times, the system undergoes spinodal decomposition followed by nucleation of domains of B, for example, of one liquid in the background of the other [3, 7].
-These domains coarsen [3] and the kinetics of coarsening begins with the conservation relation that equates the number of molecules $N_{B}$ leaving (for example) a shrinking B domain (and eventually being drawn into a growing B domain), to their flux $J_{BA}$ at the interface of the B domain with the background of A: $dN_{B}/dt=-J_{BA}S$, where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary.
+These domains coarsen [3] and the kinetics of coarsening are based upon the conservation relation that equates the number of molecules $N_{B}$ leaving (for example) a shrinking B domain (and eventually being drawn into a growing B domain), to their flux $J_{BA}$ at the interface of the B domain with the background of A: $dN_{B}/dt=-J_{BA}S$, where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary.
 For scaling purposes, we can ignore the vectors, keeping in mind that the flux and velocity of the B-molecules exiting the B-rich domain are in the direction normal to the domain boundary.
@@
 […]
-Using the conservation equation, we now write a scaling relation: $\Delta c\ \dot{\lambda}\sim-D\ \nabla\sigma/\lambda$ where $\Delta c$ is the concentration difference of the B molecules in the B-rich phase and the A-rich phase.
-$\nabla$ now is the extent of the chemical potential gradient in the direction normal to the surface of the B domain, evaluated at the domain interface.
-LSW scaling for the domain growth is obtained by approximating the gradient term by $\sigma/\lambda^{2}$ which originates in the solution of the diffusion equation for the B molecules in the background of A, at the domain surface [3].
+Using the conservation equation, we now write a scaling relation: $\Delta c\ \dot{\lambda}\sim-D\ \gamma/(\lambda\ \ell)$ where $\Delta c$ is the concentration difference of the B molecules in the B-rich phase and the A-rich phase.
+$\ell$ now is the extent of the chemical potential gradient in the direction normal to the surface of the B domain, evaluated at the domain interface.
+LSW scaling for the domain growth is obtained by approximating $\ell$ by $\lambda$ which originates in the solution of the diffusion equation for the B molecules in the background of A, at the domain surface [3] and Appendix C.
 This equation in the fast-diffusion limit (applicable to molecular diffusion in a liquid relative to the coarsening times), results in an equation for the spatial profile of the B molecule concentration in the A-rich domain where $\lambda$ determines the scale of the gradient.
 Below, we generalize the approximation for the gradient to include diffusion.
-The result for a shrinking domain gives $\dot{\lambda}\sim-(D/\Delta c)\,\sigma/\lambda^{2}$.
-For scaling, we can take $\Delta c\sim c_{B}$ (not valid near a critical point) which gives the LSW scaling, $\lambda\sim\left(\ (D\ \sigma/c_{B})\ t\right)^{1/3}$, where we do not distinguish shrinking from growing vis a vis the scaling; the two are simply related by conservation of the B molecules.
+The result for a shrinking domain gives $\dot{\lambda}\sim-(D/\Delta c)\,\gamma/\lambda^{2}$.
+For scaling, we can take $\Delta c\sim c_{B}$ (not valid near a critical point) which gives the LSW scaling, $\lambda\sim\left(\ (D\ \gamma/c_{B})\ t\right)^{1/3}$, where we do not distinguish shrinking from growing vis a vis the scaling; the two are simply related by conservation of the B molecules.
 As discussed below, the equations for the chemical potential gradients can be transformed into equations for the osmotic pressure gradients of the B molecules in the A-rich phase.
-II.2 Generalizing LSW to coarsening of solvent-rich domains in gels:
-In the following, the gel properties such as interfacial tension with a solvent-rich domain, elastic modulus $G$, mesh size $\xi$ are envisioned as those that characterize the equilibrium gel.
+II.2 Generalizing LSW to coarsening of solvent-rich domains in viscoelastic gels:
+In the following, the gel properties such as interfacial tension with a solvent-rich domain, elastic modulus $G$, mesh size $\xi$ those that characterize the equilibrium gel.
 We again begin with the conservation relation, $\dot{N}_{B}\sim-\Delta c_{s}\,S\,\dot{\lambda}=-JS$ where the flux is evaluated locally at the domain boundary and is controlled by the driving force acting at that boundary.
 $\Delta c_{s}$ is the difference between the solvent concentration in the solvent-rich domain and the gel.
-The flux is given by $\vec{J}=c_{s}\vec{v}$ where $c_{s}$ solvent concentration in the solvent-rich domain which, due to capillary pressure, ejects solvent molecules into the gel with a $\vec{v}$.
-For a more detailed discussion of this and the following see Appendix C.
-However, instead of diffusing in the polymer-rich phase, the velocity is governed by solvent flow in the deformable gel considered as a porous medium where the Darcy relation [6, 1, 2, 24] that connects the velocity to the pressure gradient: $\vec{v}=(k_{0}/\eta_{D})\nabla p$ where $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity appropriate to the solvent flow impeded by its friction with the polymer network.
+The flux is given by $\vec{J}=c_{s}\vec{v}$ where $c_{s}$ solvent concentration in the solvent-rich domain which, due to capillary pressure, releases solvent molecules into the gel with a velocity $\vec{v}$.
+For a more detailed derivation of this and the following, see Appendix C.
+However, instead of diffusing in the polymer-rich phase, the solvent flows in the deformable, viscoelastic gel considered as a porous medium.
+The flow velocity $\vec{v}$ is given by the Darcy relation [6, 1, 2, 23] that relates this velocity to the pressure gradient in the solvent pores of the gel: $\vec{v}=(k_{0}/\eta_{D})\nabla p$.
+Here $k_{0}$ is the permeability of the gel and $\eta_{D}$ is the viscosity appropriate to the solvent flow impeded by its friction with the polymer network.
 Because this transport is controlled by the hydrodynamic permeability of the solvent rather than by polymer rearrangements, the network topology is effectively fixed on the Darcy-flow timescale: polymer configurations do not relax rapidly enough to eliminate the stress induced by solvent flow.
@@
 The pressure gradient entering Darcy’s law is the local gradient at the interface of the solvent-rich domain and the gel, whose magnitude is set by how the interfacial pressure perturbation spreads into the gel.
-In incompressible systems pressure gradients are proportional to chemical potential gradients of the solvent molecules in the gel, by the thermodynamic relation, $\nabla\mu=\nabla p/c_{B}$; see Eq.
+In incompressible fluid systems pressure gradients are proportional to chemical potential gradients of the solvent molecules in the gel, by the thermodynamic relation, $\nabla\mu=\nabla p/c_{B}$; see Eq.
 1.1.14 of Ref.
 [19] and Appendix C.
-The chemical potential at the solvent-rich domain surface is again given by capillarity, $\delta\mu=\sigma/(\lambda c_{s})$ and the osmotic pressure jump at the interface (which acts as the mechanical pressure jump) is given by the Laplace law as $\delta p=c_{s}\delta\mu=\sigma/\lambda$.
+The chemical potential at the solvent-rich domain surface is again given by capillarity, $\delta\mu=\gamma/(\lambda c_{s})$ and the osmotic pressure jump at the interface (which acts as the mechanical pressure jump) is given by the Laplace law as $\delta p=c_{s}\delta\mu=\gamma/\lambda$.
 Finally, the conservation law and the Darcy relation used in the flux give the dynamical scaling relation for a shrinking domain:
-$\dot{\lambda}\sim-\frac{k_{0}\ c_{s}}{\eta_{D}}\nabla p\sim-\frac{k_{0}\ \sigma}{\eta_{D}\ \lambda\ \ell_{p}}$ (1)
+$\dot{\lambda}\sim-\frac{k_{0}\ c_{s}}{\eta_{D}}\nabla p\sim-\frac{k_{0}\ \gamma}{\eta_{D}\ \lambda\ \ell_{p}}$ (1)
 where $\ell_{p}$ is the scale of the pressure gradient, evaluated at the interface of the solvent-rich domain and the gel, and where we have set $c_{s}/\Delta c_{s}$ to be a dimensionless constant of order unity for scaling purposes.
 The solvent pressure in the gel decreases from its maximal value at the boundary of the shrinking domain via stress diffusion, known as poroelasticity [23, 4].
-That both the stress and pressure propagate diffusively can be understood from the low-Reynolds number force balance in a gel which implies that force balance implies that pressure gradients are balanced by stress gradients.
+That both the stress and pressure propagate diffusively can be understood from the low-Reynolds number force balance in a gel where pressure gradients are balanced by viscoelastic, polymer stress gradients.
 At relatively early times, when the poroelastic diffusion length $\ell_{p}$ is smaller than the domain size $\lambda$, the pressure gradient at the domain boundary can be estimated as $\nabla p\sim\delta p/\ell_{p}$; this approximation captures the transient buildup of the interfacial gradient before it reaches its long-time, saturated value.
@@
 This description applies only until $\ell_{p}$ reaches the domain size $\lambda(t)$.
-Beyond this point, the pressure field has equilibrated across the entire solvent-rich domain, and extending the gradient over larger length scales does not increase the capillary driving force acting on that domain.
-Mathematically, this is because the finite domain size, together with the Laplace pressure boundary condition at the domain interface, fix both the magnitude and the spatial extent of the pressure gradients.
+As shown in Appendix C, this is because the finite domain size, together with the Laplace pressure boundary condition at the domain interface, fix both the magnitude and the spatial extent of the pressure gradients.
 Those are substantial only for distances from the domain smaller or equal to $\lambda$.
-As shown in more detail in Appendix C, beyond this point, the solution of the poroelastic diffusion equation has evolved so that the pressure gradient evaluated at the domain interface no longer changes with time, and further spreading of the pressure field to larger distances does not increase the capillary driving force acting on that domain.
+Beyond this point, the solution of the poroelastic diffusion equation has evolved so that the pressure gradient evaluated at the domain interface no longer changes with time.
 We note that the elastic modulus enters the poroelastic diffusion constant even when the gel is in its viscoelastic regime because the modulus controls the initial magnitude of the viscoelastic stress.
@@
 We note that most treatments of poroelasticity take the crosslinked gel to have an elastic (as opposed to viscoelastic) stress response [25, 12].
-This indeed applies for times greater than the intrinsic viscoelastic crossover time of the gel, $\tau_{\mathrm{el}}$, which separates the early-time stress response, when polymer motions can relax an imposed stress, from the long-time elastic response [11, 22].
+This indeed applies for times greater than the intrinsic viscoelastic crossover time of the gel, $\tau_{\mathrm{el}}$, which separates the early-time viscoelastic stress response, from the long-time elastic response [11, 22].
 For $t\gg\tau_{\mathrm{el}}$, the shorter-time polymeric degrees of freedom have equilibrated, the crosslinked polymers behave elastically, and dissipation associated with stress redistribution is controlled primarily by solvent flow relative to a static polymer network.
 In this regime, the viscosity entering stress diffusion is well approximated by the Darcy viscosity, so that $\eta_{ps}\approx\eta_{D}$.
-During coarsening, however, the gel is responding dynamically to stresses imposed by solvent being ejected from shrinking domains and accumulated in growing domains.
+During coarsening, however, the gel is responding dynamically to stresses imposed by solvent being released from shrinking domains and accumulated in growing domains.
 These stresses act as localized perturbations to a gel that is otherwise close to its equilibrium swollen state.
 Because these perturbations are time dependent and spatially localized, they do not probe the static, long-time elastic response of the network, but instead drive transient polymer stress relaxation.
-Permanent crosslinking suppresses macroscopic flow but does not eliminate internal polymer dissipation; even dry covalently crosslinked networks exhibit viscoelastic stress relaxation at intermediate times due to strand-level and non-affine relaxation modes ($\sim$ a few hundred s), before reaching an elastic plateau [25].
+Permanent crosslinking suppresses macroscopic polymeric flow but does not eliminate internal polymer dissipation; even dry covalently crosslinked networks exhibit viscoelastic stress relaxation at intermediate times due to polymeric (individual chain and local, transient cooperative) relaxation modes , before reaching an elastic plateau [24].
 For this reason, we envision the gel responding viscoelastically on the timescale of coarsening, before elastic arrest intervenes and halts the dynamics at long times (on the scale of hours).
@@
 The local and transient polymer rearrangements contribute to stress relaxation, and the relevant dissipation expressed by the effective stress-relaxation viscosity $\eta_{ps}$ reflects both solvent–polymer friction and the intrinsic viscoelastic dynamics of the polymer-rich phase [22, 18].
-Accordingly, we take the viscosity entering poroelastic diffusion during coarsening to be a poroelastic viscosity $\eta_{ps}$ that is distinct from the Darcy viscosity $\eta_{D}$.
-The Darcy viscosity characterizes solvent flow in response to pressure gradients in a static network, whereas $\eta_{ps}$ governs the coupled relaxation of pressure and polymer stress when the network itself is dynamically responding.
+Accordingly, we take the viscosity entering poroelastic diffusion during coarsening to be a poroelastic viscosity $\eta_{ps}$ that is distinct from the Darcy viscosity $\eta_{D}$ which characterizes solvent flow in response to pressure gradients in a static network.
+In contrast, $\eta_{ps}$ governs the coupled relaxation of pressure and polymer stress when the network itself is dynamically responding.
 This distinction must be made even in a permanently crosslinked gel.
-There too solvent transport generates strains on timescales comparable to polymer stress relaxation.
+There too, solvent transport generates strains on timescales comparable to polymer stress relaxation.
 In general, therefore, $\eta_{ps}\neq\eta_{D}$ for $t\ll\tau_{\mathrm{el}}$.
@@
 When poroelastic diffusion governs the relaxation of pressure gradients, Eq. 1 can be written in the scaling form
-$\dot{\lambda}\sim-\frac{k_{0}\sigma}{\eta_{D}\,\lambda\,\ell_{p}},$ (3)
+$\dot{\lambda}\sim-\frac{k_{0}\gamma}{\eta_{D}\,\lambda\,\ell_{p}},$ (3)
 where $\ell_{p}$ denotes the characteristic length over which pressure relaxes in the gel.
@@
 3 yields the scaling law
-$\lambda\sim\left(\frac{k_{0}\,\sigma}{\eta_{D}\,D_{p}^{1/2}}\right)^{1/2}t^{1/4}.$ (4)
+$\lambda\sim\left(\frac{k_{0}\,\gamma}{\eta_{D}\,D_{p}^{1/2}}\right)^{1/2}t^{1/4}.$ (4)
 As above, shrinking and growing domains obey the same scaling by solvent conservation.
 In this slow-poroelastic regime, the coarsening law exhibits an explicit dependence on the gel modulus through the poroelastic diffusion constant, $D_{p}\sim G$, giving $\lambda\sim G^{-1/4}t^{1/4}$.
-The extent to which the prefactor $\sqrt{k_{0}\sigma/\eta_{D}}$ introduces additional modulus dependence is discussed further in the Results.
-The use of the diffusion length $\ell_{p}=\sqrt{D_{p}\ t}$ for the slow poroelastic limit is only appropriate if $\ell_{p}<\lambda(t)$; relaxation of the gel on scales larger than the domain size $\lambda(t)$ is not relevant for solvent injection into the gel by a shrinking domain of size $\lambda$.
+The extent to which the prefactor $\sqrt{k_{0}\gamma/\eta_{D}}$ introduces additional modulus dependence is discussed further in the Results.
+The use of the diffusion length $\ell_{p}=\sqrt{D_{p}\ t}$ for the slow poroelastic limit is only appropriate if $\ell_{p}<\lambda(t)$; relaxation of the gel on scales larger than the domain size $\lambda(t)$ is not relevant for solvent release n into the gel by a shrinking domain of size $\lambda$ (see Appendix C).
 Thus, the slow poroelastic limit crosses over to the fast limit at a crossover domain size.
 Equating $\ell_{p}\sim\lambda(t)$ yields a time-independent crossover length
-$\lambda_{\times}\sim\,\frac{\sigma}{G}\,\frac{\eta_{ps}}{\eta_{D}},$ (5)
-which magnifies the elastocapillary length $\sigma/G$ (which is of the order the mesh size of the gel) by a viscosity ratio.
+$\lambda_{\times}\sim\,\frac{\gamma}{G}\,\frac{\eta_{ps}}{\eta_{D}},$ (5)
+which magnifies the elastocapillary length $\gamma/G$ (which is of the order the mesh size of the gel) by a viscosity ratio.
 The crossover length can be large since the Darcy viscosity $\eta_{D}$ accounts for the solvent viscosity and friction with the instantaneous state of the polymers of the gel while the poroelastic viscosity in the viscoelastic regime of the gel, where coarsening takes place, accounts for polymer-polymer friction and local chain relaxation When the domain size $\lambda(t)<\lambda_{\times}$ the slow limit applies with coarsening scaling as $t^{1/4}$.
 In the opposite limit, the pressure gradients decay as $1/\lambda(t)$ which is appropriate in the fast poroelastic limit where the domain size scaling is $t^{1/3}$.
-II.3 Elastic arrest of coarsening
+III Elastic arrest
 So far, we have focused on the capillary driving force for coarsening.
-At early times, the coarsening gel is in the viscoelastic regime, where polymers respond locally and transiently to the stress imposed by solvent expelled from shrinking domains.
-In this regime, elastic stresses relax on times shorter than the coarsening time and there is no spatially long-ranged and temporally persistent elastic response; those transient stresses therefore act primarily as a source of dissipation, slowing—but not halting—the capillarity-driven coarsening dynamics.
-At sufficiently long times, $t>\tau_{\mathrm{el}}$, the local and transient viscoelastic polymer relaxation modes coupled to the coarsening-induced stress have relaxed, and the polymer network responds collectively and elastically [11, 22].
-Although the elastic modulus is a property of the network at all times, it becomes dynamically operative as a long-ranged elastic stress that resists solvent flow only in the elastic regime, when stresses generated by solvent transport can no longer relax locally and a temporally permanent and a spatially long-ranged stress profile is established in the gel.
-In an elastic gel, capillary forces are balanced by gel stress at the domain boundary, determined self-consistently by elastic deformation of the entire surrounding medium.
-Elastic arrest is therefore most transparently understood as a force balance at the boundary of the shrinking domain: the Laplace capillary pressure ($\delta p=\sigma/\lambda)$ is canceled by the elastic stress of the gel.
-Crucially, the elastic stress at the domain boundary is not fixed a-priori.
+At early times, the coarsening gel is in the viscoelastic regime, where polymers respond locally and transiently to the stresses generated by capillary tractions (normal stresses) at curved solvent–gel interfaces.
+In this regime, there is no spatially long-ranged and temporally persistent elastic response; transient polymer stresses therefore act primarily as a source of dissipation, slowing—but not halting—the capillarity-driven coarsening dynamics.
+Mechanically, the Young–Laplace condition prescribes a normal traction (normal stress) at the domain boundary; because a long-ranged, persistent deviatoric stress field is not possible in the viscoelastic regime, this traction is transmitted primarily through an isotropic pressure perturbation in the solvent pores (i.e., the pore-pressure) of the gel.
+Matching this perturbation to the far-field pressure necessarily generates a nonzero pressure gradient, which by Darcy’s law drives solvent flow and hence coarsening.
+However, at sufficiently long times, $t>\tau_{\mathrm{el}}$, the local and transient viscoelastic polymer relaxation modes coupled to the coarsening-induced stress have relaxed, and the polymer network responds collectively and elastically [11, 22].
+Although the elastic modulus is a property of the network at all times, the gel stress is only spatially long-ranged and temporally persistent in this elastic regime.
+In this regime, the Young–Laplace interfacial traction (i.e., the normal force per unit area—equivalently, the normal component of the stress—transmitted across the interface) has a characteristic magnitude set by the Laplace-pressure scale, $\delta p\sim\gamma/\lambda$, at the solvent–gel interface.
+As shown in Appendix D, boundary traction is then balanced by the long-ranged deviatoric (non-isotropic) stress field established in the gel, but the pressure gradient is zero.
+Crucially, the resulting pore-pressure field is not assumed to have zero gradient: it is determined by solving the elastic equilibrium problem in the surrounding gel.
+Writing the gel stress as $\bm{\sigma}=-p\,\mathbf{I}+2G\bm{\varepsilon}$ and imposing spherical mechanical equilibrium, the elastic displacement/strain field that balances the interfacial traction yields $dp/dr=0$, i.e.
+the isotropic pore pressure outside the domain is spatially uniform (with the constant fixed by the far-field boundary condition) (see Appendix D).
+Therefore, in this elastic regime the pressure gradient entering Darcy’s law vanishes, so the pressure-gradient mechanism for solvent transport is eliminated.
+By Darcy’s law, this removes the driving force for solvent flow and thus halts the capillarity-driven coarsening.
+We therefore refer to elastic arrest as the regime in which the interfacial capillary traction is balanced by the gel’s elastic shear stress, leaving no pore-pressure gradient to drive further Darcy permeation.
+It is also important to note that the elastic stress at the domain boundary is not fixed a-priori.
 It must be determined by solving the elastic displacement problem in the gel surrounding the entire domain.
-For a spherical domain in an elastic solid [14], normal interfacial stress is of order
-$\sigma_{\rm el}\;\sim\;G\,\frac{\delta\lambda}{\lambda}$ (6)
-where $G$ is the shear modulus of the gel and $\delta\lambda$ is the elastic displacement of the interface.
-Physically, $\delta\lambda$ represents the instantaneous radial deformation of the gel at the domain boundary induced by the capillary pressure acting on the curved interface; in the elastic regime it is defined at all times as the solution of the elastic equilibrium equations with the instantaneous interfacial pressure $\delta p(t)\sim\sigma/\lambda(t)$ applied as a boundary condition, which in general is time dependent because $\lambda(t)$ evolves during coarsening.
-Prior to arrest, the magnitude of $\delta\lambda$ is smaller than its long-time value because the long-ranged elastic stress field has not yet fully developed and therefore lags behind the capillary pressure.
+For a spherical domain embedded in an infinite elastic gel [14], the radial displacement field decays as (see Appendix D)
+$u(r)\sim\delta\lambda\left(\frac{\lambda}{r}\right)^{2},$ (6)
+so that the normal stress at the interface is proportional to the interfacial strain,
+$\sigma_{\rm el}\sim G\frac{\delta\lambda}{\lambda},$ (7)
+up to a numerical prefactor.
+Physically, $\delta\lambda$ represents the instantaneous radial deformation of the gel at the domain boundary induced by the capillary pressure acting on the curved interface.
+It is not an imposed equilibrium mismatch between two stress-bearing phases; rather, it arises dynamically from the requirement that the gel satisfy mechanical equilibrium with the instantaneous Laplace pressure $\delta p(t)\sim\gamma/\lambda(t)$.
+In the elastic regime it is defined at all times as the solution of the elastic equilibrium equations with the instantaneous interfacial pressure applied as a boundary condition, which in general is time dependent because $\lambda(t)$ evolves during coarsening.
 At long times, once the elastic stress has equilibrated spatially, the Laplace force balance at the domain boundary fixes $\delta\lambda$ by equating the elastic stress to the capillary pressure, $G(\delta\lambda/\lambda)\sim\gamma/\lambda$, which yields $\delta\lambda\sim\gamma/G$; thus, although the domain size $\lambda$ is mesoscopic, the interfacial displacement itself is set by a microscopic elastic length of order the gel mesh size.
-Because the elastic deformation extends over a region of size $\sim\lambda$, the domain size itself enters the solution: larger domains require elastic deformation over larger distances and therefore generate a different stress response.
-By contrast, the elastic stress at the solvent-rich domain boundary is not present in the viscoelastic, coarsening regime, where the unbalanced capillary pressure $\delta p$ is what drives solvent flow and coarsening.
+Importantly, this boundary force balance fixes the elastic displacement $\delta\lambda$ but it does not by itself determine the arrested domain size itself $\lambda_{\rm arrest}$ which is determined kinetically as discussed below.
+To summarize: In the elastic regime of the gel there is only a deviatoric (non-isotropic) stress induced by the capillary traction at the boundary of the solvent-rich domain and the pressure gradient is zero (see Appendix D).
+Since $\dot{\lambda}\sim-(k_{0}/\eta_{D})\nabla p$, the solvent flux ceases and coarsening stops.
+By contrast, in the viscoelastic coarsening regime, the long-ranged elastic stress field has not yet developed, so the capillary pressure remains effectively unbalanced, a pressure gradient exists in the gel and by the Darcy flow equation solvent flow and coarsening proceed.
+Such elastic arrest does not occur in a coherent AABB solid solution (of two elastic phases), since mass transport is diffusive and driven by gradients of chemical potential, which is not proportional to the pressure gradient in an elastic system.
+The chemical potential contain contributions from the entire elastic stress field (both isotropic pressure and deviatoric, non-isotropic stress).
+Consequently, in the gel elastic arrest of the solvent-rich, fluid domains occurs when $\nabla p=0$, whereas in a solid solution elasticity modifies the thermodynamic driving force itself rather than eliminating transport through the vanishing of a pressure gradient.
 We emphasize that elastic arrest is a kinetic effect that is relevant only for $t>\tau_{\mathrm{el}}$.
@@
 Using the slow-poroelastic time-to-size relation of Eq. 4 ($\lambda^{4}\propto t$), we obtain the arrest size in the slow-diffusion regime,
-$\lambda_{\rm arrest}^{(\rm slow)}\sim\Bigg(\,\frac{\eta_{ps}\,\eta_{el}}{\eta_{D}^{2}}\,\frac{k_{0}\,\sigma^{2}}{G^{2}}\,\Bigg)^{1/4},$ (7)
+$\lambda_{\rm arrest}^{(\rm slow)}\sim\Bigg(\,\frac{\eta_{ps}\,\eta_{el}}{\eta_{D}^{2}}\,\frac{k_{0}\,\gamma^{2}}{G^{2}}\,\Bigg)^{1/4},$ (8)
 where we have expressed the elastic crossover time in terms of an effective elastic viscosity $\eta_{el}=G\tau_{\mathrm{el}}$, because viscosities scale as a modulus times a characteristic relaxation time.
@@
 In the fast-poroelastic regime, the corresponding arrest size is
-$\lambda_{\rm arrest}^{(\rm fast)}\sim\Bigg(\,\frac{\eta_{el}}{\eta_{D}}\,\frac{k_{0}\,\sigma}{G}\,\Bigg)^{1/3}.$ (8)
-III Results
+$\lambda_{\rm arrest}^{(\rm fast)}\sim\Bigg(\,\frac{\eta_{el}}{\eta_{D}}\,\frac{k_{0}\,\gamma}{G}\,\Bigg)^{1/3}.$ (9)
+IV Results
 In this section, we analyze the previous results to predict the scaling of the domain coarsening as a function of time and gel elastic modulus, as well as the modulus dependence of the arrest length.
 In addition, we show that this length can be magnified to a value far larger than the gel mesh size, by the viscosity ratios defined above.
-We present our analysis for two limiting cases of the dependence of the gel properties (permeability $k_{0}$ and interfacial tension with the solvent-rich phase $\sigma$) that enter the capillary-driven coarsening and eventual arrest.
-We keep the explicit dependence on the gel modulus $G$ since this is often changed and measured in experiments [9].
+We present our analysis for two limiting cases of the dependence of the gel properties (permeability $k_{0}$ and interfacial tension with the solvent-rich phase $\gamma$) that enter the capillary-driven coarsening and eventual arrest.
+We keep the explicit dependence on the gel modulus $G$ since this is often varied and measured in experiments [9].
 The first scenario we call the mesh-scaling gel where the polymer fraction in the gel is relatively small and the mesh size $\xi$ is the largest length in the homogeneous gel.
@@
 This is the limit that is applicable to the gels in [9] where the arrest length has been measured.
-For the melt-like gel, we keep $k_{0}$ and $\sigma$ as parameters which vary slowly with the gel mesh size, due to the involvement of the other length scales of the system.
-We first present the scaling relations for the coarsening dynamics and the arrest length as functions of the three viscosities defined above: The Darcy viscosity $\eta_{D}$, the poroelastic viscosity for stress diffusion in the viscoelastic regime of the polymers $\eta_{ps}$, and the viscosity $\eta_{el}=G\ \tau_{el}$ related to the time scale $\tau_{el}$, at which the viscoelastic regime crosses over to the elastic regime.
+For the melt-like gel, we keep $k_{0}$ and $\gamma$ as parameters which vary slowly with the gel mesh size, due to the involvement of the other length scales of the system.
+We first present the scaling relations for the coarsening dynamics and the arrest length as functions of the three viscosities defined above: The Darcy viscosity $\eta_{D}$, the poroelastic viscosity for stress diffusion in the viscoelastic regime of the polymers $\eta_{ps}$, and the viscosity $\eta_{el}=G\ \tau_{el}$ related to the time scale $\tau_{el}$, at which the viscoelastic regime (present at short times even in a crosslinked gel) crosses over to the elastic regime.
 Afterwards, we rewrite the scaling relations as functions of time scales associated with the Darcy and poroelastic viscosities, guided by the fact that a viscosity is the product of a modulus and a time scale.
@@
 Gel Model Regime Coarsening law $\lambda(t)$ Arrest length $\lambda_{\rm arrest}$
-(i) Mesh-scaling gel: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim\left(\frac{k_{B}T}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
+(i) Mesh-scaling gel: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\gamma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim\left(\frac{k_{B}T}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
 Slow $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/12}\left(\frac{\eta_{ps}}{\eta_{D}^{2}}\right)^{1/4}t^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
-(ii) Melt-like gel: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
-Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G}\,\frac{\eta_{ps}}{\eta_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
+(ii) Melt-like gel: $\gamma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\gamma}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\gamma}{G}\,\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
+Slow $\lambda(t)\sim\left(\frac{k_{0}\gamma^{2}}{G}\,\frac{\eta_{ps}}{\eta_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
 Table 2: Same as Table 1, but rewriting all viscosities as products of an elastic modulus and characteristic time scales.
 Gel Model Regime Coarsening law $\lambda(t)$ Arrest length $\lambda_{\rm arrest}$
-(i) Mesh-scaling gel + $\eta\sim G\tau$: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
+(i) Mesh-scaling gel + $\eta\sim G\tau$: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\gamma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
 Slow $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
-(ii) Melt-like gel + $\eta\sim G\tau$: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{G}\,\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
-Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
+(ii) Melt-like gel + $\eta\sim G\tau$: $\gamma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\gamma}{G}\,\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\gamma}{G}\,\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
+Slow $\lambda(t)\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
 We note that in Table 2 where we have extracted the maximal dependence on the elastic modulus $G$ that is physically reasonable, the fast and slow coarsening for the mesh-scaling gel both scale as $1/{G}^{1/3}$ with respective time dependence of $t^{1/3}$ and $t^{1/4}$ respectively.
@@
 This is also the modulus scaling of the arrest length in the slow poroelastic limit.
-IV Discussion
+V Discussion
 Interpretation of mesh-scaling gel: scaling of domain coarsening with time and elastic modulus: The only length scale in the problem is the gel mesh size $\xi\sim 1/G^{1/3}$.
@@
 Nevertheless, the “fast” and “slow” nomenclature refers to the coarsening exponents rather than the characteristic time scales: even when the slow poroelastic mode is governed by a shorter characteristic time scale, it still yields slower growth in time ($t^{1/4}$ versus $t^{1/3}$).
-Melt-scaling gel: scaling of domain coarsening with time and elastic modulus: For this system, we keep the porosity and interfacial tension as parameters that vary only slowly with the modulus $G$, since for large polymer fraction gels (the melt-scaling limit), additional length scales related to the Kuhn length and solvent size can determine these parameters.
-However, despite this, the crosslinked, melt-scaling gel is still well characterized by its elastic modulus $G$ which remains the dominant parameter controlling both stress transmission and arrest.
+Melt-like gel: scaling of domain coarsening with time and elastic modulus: For this system, we keep the porosity and interfacial tension as parameters that vary only slowly with the modulus $G$, since for large polymer fraction gels (the melt-like limit), additional length scales related to the Kuhn length and solvent size can determine these parameters.
+However, despite this, the crosslinked, melt-like gel is still well characterized by its elastic modulus $G$ which remains the dominant parameter controlling both stress transmission and arrest.
 In this regime, fast poroelastic diffusion predicts coarsening which scales as $t^{1/3}$ while slow poroelastic diffusion coarsening scales as $t^{1/4}$.
@@
 The arrest length is magnified from its microscopic or mesh-size value by the ratio of the time scales $\tau_{el}\gg\tau_{D}$ for fast coarsening and $\tau_{el}\tau_{ps}\gg\tau_{D}^{2}$ for slow coarsening.
-These large ratios reflect the separation between solvent transport and polymer stress relaxation processes: $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains activated by the solvent flow while $\tau_{D}$ is related to solvent flow in the presence of polymer friction in an essentially static configuration.
+These large ratios reflect the separation between solvent transport and polymer stress relaxation processes: $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains activated by the solvent flow, while $\tau_{D}$ is related to solvent flow in the presence of polymer friction in an essentially static configuration.
 The longest time scale is $\tau_{el}$ which is the crossover from viscoelasticity (activated by solvent flow) to elastic response when further coarsening is kinetically arrested by elastic stress acting over a region of order $\lambda$.
@@
 Our scaling relations also predict the dependence of the arrest-length on the various viscosities and future measurements of this behavior (see below) would be of interest.
-We now address the very slow drainage of solvent at time scales of about a day.
-For that process, also driven by capillarity involving the solvent in the gel and the excess solvent reservoir, the solvent flow is slow and we expect that the macroscopic drainage to depend on poroelastic diffusion, but with Darcy flow related to the solvent viscosity $\eta_{s}$ and not $\eta_{ps}$ since the polymers have relaxed all stresses on the macroscopic time scale.
-We thus estimate that the macroscopic time for drainage scales as $R^{2}/D_{D}$ where $R$ is the macroscopic sample size and the Darcy poroelastic diffusion constant $D_{D}$ scales as $k_{0}G/\eta_{s}$.
+We now address the very slow drainage of solvent at time scales of many hours.
+That process, is also driven by capillarity involving the solvent in the gel near its boundary with an excess solvent reservoir.
+The solvent flow is slow and we expect that the drainage initiated at the sample boundaries (not included in the other sections of this paper) will depend on poroelastic diffusion in that region, with the Darcy flow near the boundary related to the solvent viscosity $\eta_{s}$ and not $\eta_{ps}$; this is because the polymers have relaxed all stresses on the macroscopic time scale.
+We thus estimate that the macroscopic time for drainage at the boundary scales as $R^{2}/D_{D}$ where $R$ is a macroscopic size and the Darcy poroelastic diffusion constant $D_{D}$ scales as $k_{0}G/\eta_{s}$.
 For an order of magnitude numerical estimate, we take the permeability $k_{0}\sim\xi^{2}\sim(k_{B}\ T/G)^{2/3}$ and $\eta_{S}$ to be of the order of magnitude of the water viscosity, $10^{-3}\text{Pa~s}$.
@@
 Such measurements would allow a direct comparison between macroscopic poroelastic response and microscopic coarsening dynamics.
-V Acknowledgements
+VI Acknowledgements
 We are grateful to Eric Dufresne and Carla Fernandez-Rico for discussions about their experiments and to David Andelman, Haim Diamant, Amit Kumar, Biplab Mandal, Manu Mannattil, and David Zwicker for insights into their theoretical models.
-Appendix A Onuki-Puri numerical calculations
+Appendix A Curk and Luijten model
+A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\gamma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix.
+Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects, and does not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach (see Appendix for further discussion).
+They predict an initial regime of growth followed by elastic arrest at short times, an intermediate viscoelastic coarsening regime with linear-in-time growth, and a late-time crossover back to LSW coarsening with domain size scaling as $t^{1/3}$.
+The late-time return to LSW scaling is not observed in PDMS gels, where the long-time (hours) state is one of elastic arrest, indicating a distinct rate-limiting physical mechanism.
+Appendix B Onuki-Puri numerical calculations
 A previous, numerical calculation by Onuki and Puri [17] focused on a two-dimensional crosslinked gel with (implicit) solvent-gel friction.
@@
 In their simulations, the effective network–solvent friction includes an explicit Stokes prefactor ($6\pi$), which enters directly into the arrest criterion.
-As a result, the arrested length can be enhanced by roughly an order of magnitude relative to the naive elastocapillary estimate $\sigma/G$, even in the absence of an additional large control parameter.
+As a result, the arrested length can be enhanced by roughly an order of magnitude relative to the naive elastocapillary estimate $\gamma/G$, even in the absence of an additional large control parameter.
 By contrast, in our scaling theory the viscosities enter only through ratios (see the Tables below), so such numerical prefactors cancel.
 The remaining microscopic input enters the arrest length only to a fractional power, rendering the predicted mesoscopic magnification robust and weakly sensitive to microscopic details.
-Appendix B Curk and Luijten model
-A paper with theory and model calculations of phase separation and coarsening in a viscoelastic gel by Curk and Luijten [5] extended LSW-type Ostwald ripening driven by capillarity (interfacial tension $\sigma$) by incorporating elastic stresses and their viscoelastic relaxation in the surrounding matrix.
-Their approach is based on diffusion-controlled evaporation–condensation through a viscoelastic solid, modified by elastic and viscoelastic effects, and does not include poroelastic (Darcy) solvent flow through a deformable, porous network, which provides the dominant dissipative mechanism for relaxing capillary pressure in our approach (see Appendix for further discussion).
-They predict an initial regime of growth followed by elastic arrest at short times, an intermediate viscoelastic coarsening regime with linear-in-time growth, and a late-time crossover back to LSW coarsening with domain size scaling as $t^{1/3}$.
-The late-time return to LSW scaling is not observed in PDMS gels, where the long-time (hours) state is one of elastic arrest, indicating a distinct rate-limiting physical mechanism.
-Appendix C Unified diffusion-based description of coarsening in fluids and gels
+Appendix C Unified diffusion-based description of coarsening in fluids and viscoelastic gels
 In this Appendix we present a unified derivation of the coarsening dynamics for (i) phase separation in a binary fluid mixture of small molecules, and (ii) solvent-rich domains embedded in a polymer gel.
@@
 where $c_{B}$ is the number density of the transported species.
-Equivalently, using $\nabla\mu=\nabla p/c_{B}$, the flux may be written as $J\sim-D\nabla c_{B}$.
+Equivalently, using $\nabla\mu=\nabla p/c_{B}$ which is appropriate for fluids, the flux may be written as $J\sim-D\nabla c_{B}$.
 In this case, the mobility is due to the rapid diffusive motion of individual molecules in the surrounding fluid.
@@
 Thus, while the diffusion equation for the pressure field has the same mathematical form in both systems, the mobility MM—and therefore the physical origin of dissipation—is fundamentally different in binary fluids and in gels.
-The capillary driving force originates in the curvature dependence of the chemical potential at the interface.
+In both cases the capillary driving force originates in the curvature dependence of the chemical potential at the interface.
 For a spherical domain of radius $\lambda$, the chemical potential at the interface exceeds that of a flat interface by
-$\delta\mu=\frac{\sigma}{c\,\lambda},$ (17)
-where $\sigma$ is the interfacial tension and $c$ is the concentration of the transported species in the domain.
+$\delta\mu=\frac{\gamma}{c\,\lambda},$ (18)
+where $\gamma$ is the interfacial tension and $c$ is the concentration of the transported species in the domain.
 C.3 Relation of mechanical and osmotic pressures
@@
 In this elastic regime, further solvent transport requires deformation of the entire domain, and the resulting pressure gradients are determined by force balance across the domain rather than by local osmotic gradients alone.
-This buildup of elastic stress leads to elastic arrest of coarsening.
-Thus, while elastic stresses are essential for determining the arrest of coarsening at long times, on the coarsening time scale itself the polymer network behaves viscoelastically, and elastic stresses relax locally and do not contribute significantly to the pressure gradients that drive flow.
-Capillary stresses enter separately as boundary conditions at interfaces: the Laplace law fixes the pressure jump across the interface between the solvent-rich domain and the gel and thereby sets the magnitude of the pressure difference that must be accommodated by the pressure field in the surrounding material.
-In summary, while the mechanical pressure field includes osmotic, elastic, and capillary contributions, gradients of chemical potential may be equivalently expressed in terms of pressure gradients because osmotic stresses provide the fundamental thermodynamic driving force, with transient and local elastic stresses controlling how this driving force is transmitted in the viscoelastic regime of the gel.
+This buildup of elastic stress leads to elastic arrest of coarsening (see Appendix D.
+Thus, while elastic stresses are essential for determining the arrest of coarsening, on the coarsening time scale itself the polymer network behaves viscoelastically, and elastic stresses relax locally and do not contribute significantly to the pressure gradients that drive flow.
+Capillary stresses enter separately as boundary conditions at interfaces: the Laplace law fixes the pressure jump across the interface between the solvent-rich domain and the gel and thereby sets the magnitude of the pressure difference that is the boundary condition on the pressure field in the surrounding material.
+In summary, while the mechanical pressure field in fluids includes osmotic and capillary contributions, gradients of chemical potential may be equivalently expressed in terms of pressure gradients because osmotic stresses provide the fundamental thermodynamic driving force, with transient and local elastic stresses controlling how this driving force is transmitted (via energy dissipation) in the viscoelastic regime of the gel.
 C.4 Relation between chemical potential and pressure gradients.
@@
 Identifying the chemical potential as $\mu=df/dc$, this can be written as
-$\Pi=\mu\,c_{B}-f(c_{B}).$
-To make this relation explicit, we differentiate
-$\Pi=\mu\,c-f(c)$
-with respect to the state variables at fixed temperature. Using the chain rule, this gives
+$\Pi=\mu\,c-f(c).$
+To make this relation explicit, we differentiate $\Pi$ at fixed temperature. Using the chain rule,
 $d\Pi=\mu\,dc+c\,d\mu-df.$
@@
 $df=\mu\,dc.$
-Substituting this expression into the equation above, the terms proportional to $dc$ cancel, leaving
+Substituting, the terms proportional to $dc$ cancel, leaving
 $d\Pi=c\,d\mu.$
-This shows that gradients in chemical potential are equivalently expressed as gradients in osmotic pressure, which provide the mechanical driving force for transport. Therefore, at fixed temperature,
+Therefore, when the free energy depends only on concentration,
 $\nabla\Pi=c\,\nabla\mu.$ (20)
-We recall that $c$ is the concentration of the transported species (the B molecules in the AB small-molecule mixture or the solvent molecules in the solvent-gel system) outside the coarsening solvent-rich (or B molecule rich) domain.
+Limitation to fluid phases. The proportionality above relies on the assumption that
+$f=f(c)$
+is a function of concentration alone. In an elastic solid or gel, however, the free-energy density depends on additional independent variables such as the strain $\varepsilon$:
+$f=f(c,\varepsilon).$
+In that case,
+$df=\left(\frac{\partial f}{\partial c}\right)_{\varepsilon}dc+\left(\frac{\partial f}{\partial\varepsilon}\right)_{c}d\varepsilon,$
+and the chemical potential is defined at fixed strain,
+$\mu=\left(\frac{\partial f}{\partial c}\right)_{\varepsilon}.$
+Repeating the derivation gives
+$d\Pi=c\,d\mu-\left(\frac{\partial f}{\partial\varepsilon}\right)_{c}d\varepsilon.$
+Thus, in an elastic medium where strain varies spatially, gradients in chemical potential are not simply proportional to gradients in pressure.
+The additional mechanical term couples composition and deformation, so transport cannot in general be written purely in terms of $\nabla\Pi$.
+The relation $\nabla\Pi=c\,\nabla\mu$ holds only for systems whose thermodynamics is fully described by concentration alone (e.g.
+two-component fluids), but not for elastic systems in which strain is an independent degree of freedom.
+In the relation of the osmotic pressure and the chemical potential, $c$ denotes the concentration of the transported species (the B molecules in the AB small-molecule mixture or the solvent molecules in the solvent–gel system) outside the coarsening solvent-rich (or B-rich) domain.
 In an incompressible system, the mechanical pressure (outside the solvent-rich domain) that drives transport differs from the osmotic pressure only by a reference constant (and by elastic or capillary contributions treated separately through force balance and (Laplace) boundary conditions at the solvent-rich domain interface).
@@
 The capillary boundary condition fixes the mechanical pressure at the domain surface through the Laplace (osmotic) pressure jump $\delta p$,
-$p(r=\lambda,t)=\delta p\equiv\frac{\sigma}{\lambda}.$ (22)
-C.5 Diffusion equation for the pressure field
+$p(r=\lambda,t)=\delta p\equiv\frac{\gamma}{\lambda}.$ (23)
+C.5 Diffusion equation for the pressure field outside the coarsening domain
 We show now that the mechanical pressure field satisfies a diffusion equation,
@@
 $\left.\frac{\partial p}{\partial r}\right|_{r=\lambda}=-\delta p\left(\frac{1}{\lambda}+\frac{1}{\sqrt{\pi Dt}}\right).$ (35)
-This expression shows explicitly how the interfacial pressure gradient evolves in time as the pressure field spreads into the surrounding medium.
+where $\delta p\sim\gamma/\lambda$. This expression shows explicitly how the interfacial pressure gradient evolves in time as the pressure field spreads into the surrounding medium.
 C.7 Binary fluid: fast molecular diffusion (LSW limit)
@@
 As shown above, the pressure field outside a shrinking domain satisfies a diffusion equation and, because molecular diffusion is fast compared to the coarsening dynamics, the pressure profile rapidly reaches its long-time limit.
-Equivalently, the pressure satisfies Laplace’s equation outside the domain, with the capillary boundary condition $p(r=\lambda)=\delta p\sim\sigma/\lambda$.
+Equivalently, the pressure satisfies Laplace’s equation outside the domain, with the capillary boundary condition $p(r=\lambda)=\delta p\sim\gamma/\lambda$.
 In this limit, the interfacial pressure gradient is time independent and scales as
-$\left.\nabla p\right|_{r=\lambda}\sim\frac{\delta p}{\lambda}\sim\frac{\sigma}{\lambda^{2}}.$
+$\left.\nabla p\right|_{r=\lambda}\sim\frac{\delta p}{\lambda}\sim\frac{\gamma}{\lambda^{2}}.$
 Using the interfacial evolution law $\Delta c\,\dot{\lambda}\sim-(M/c)\left.\nabla p\right|_{r=\lambda}$ with $M\sim Dc$ for molecular diffusion then gives
-$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c}\left.\nabla p\right|_{r=\lambda}\sim-\frac{1}{\Delta c}\,D\,\frac{\sigma}{\lambda^{2}}.$ (39)
-Integrating this equation yields the standard LSW coarsening law $\lambda(t)\sim(D\sigma\,t/\Delta c)^{1/3}$.
-C.9.2 Gel (poroelastic; slow poroelastic diffusion, $\ell_{p}\ll\lambda$).
-The interfacial pressure is set by capillarity, $\delta p\sim\sigma/\lambda$, and Darcy flow gives $M\sim k_{0}c_{s}/\eta_{D}$ with $\nabla\mu=(1/c_{s})\nabla p$.
+$\dot{\lambda}\sim-\frac{1}{\Delta c}\,\frac{M}{c}\left.\nabla p\right|_{r=\lambda}\sim-\frac{1}{\Delta c}\,D\,\frac{\gamma}{\lambda^{2}}.$ (40)
+Integrating this equation yields the standard LSW coarsening law $\lambda(t)\sim(D\,\gamma\,t/\Delta c)^{1/3}$.
+C.9.2 Gel (poroelastic; slow poroelastic diffusion)
+The interfacial pressure is set by capillarity, $\delta p\sim\gamma/\lambda$, and Darcy flow gives $M\sim k_{0}c_{s}/\eta_{D}$ with $\nabla\mu=(1/c_{s})\nabla p$.
 Heree, $c_{s}$ is the solvent concentration in the gel where it is transported by coarsening of the solvent-rich domain.
-For slow poroelastic spreading, the exact diffusion solution yields $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}\sim(\sigma/\lambda)/\ell_{p}$, so that
-$\dot{\lambda}\sim-\frac{1}{\Delta c_{s}}\,\frac{k_{0}c_{s}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\frac{c_{s}}{\Delta c_{s}}\,\frac{\sigma}{\lambda\,\ell_{p}(t)},\qquad\ell_{p}(t)=\sqrt{D_{p}t},\ \ D_{p}\sim\frac{k_{0}G}{\eta_{ps}}.$ (40)
+For slow poroelastic spreading, the exact diffusion solution yields $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}\sim(\gamma/\lambda)/\ell_{p}$, so that
+$\dot{\lambda}\sim-\frac{1}{\Delta c_{s}}\,\frac{k_{0}c_{s}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\frac{c_{s}}{\Delta c_{s}}\,\frac{\gamma}{\lambda\,\ell_{p}(t)},\qquad\ell_{p}(t)=\sqrt{D_{p}t},\ \ D_{p}\sim\frac{k_{0}G}{\eta_{ps}}.$ (41)
 For scaling purposes, the dimensionless ratio $c_{s}/\Delta c_{s}$ is treated as a constant of order unity and we omit its numerical value here.
-C.9.3 Gel (poroelastic; fast poroelastic diffusion, $\ell_{p}\gg\lambda$).
-Once poroelastic diffusion has spread over distances comparable to the domain size, the interfacial gradient saturates to $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda\sim\sigma/\lambda^{2}$, and the evolution law becomes
-$\dot{\lambda}\sim-\,\frac{k_{0}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\,\frac{\sigma}{\lambda^{2}}.$ (41)
+C.9.3 Gel (poroelastic; fast poroelastic diffusion)
+Once poroelastic diffusion has spread over distances comparable to the domain size, the interfacial gradient saturates to $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda\sim\gamma/\lambda^{2}$ (see Eq.
+35), and the evolution law becomes
+$\dot{\lambda}\sim-\,\frac{k_{0}}{\eta_{D}}\left.\nabla p\right|_{r=\lambda}\sim-\frac{k_{0}}{\eta_{D}}\,\,\frac{\gamma}{\lambda^{2}}.$ (42)
 This has the same $\lambda^{-2}$ structure as LSW, but with the molecular mobility $D$ replaced by the Darcy permeation mobility $k_{0}c_{s}/\eta_{D}$, and with the additional dimensionless factor $c_{s}/\Delta c_{s}$ which we omitted for scaling purposes.
-Depending on whether the interfacial pressure gradient has reached its saturated value $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda$ or is still building up in time as $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}(t)$, one recovers either the standard LSW coarsening law or the poroelastic coarsening laws discussed in the main text.
+Depending on whether the interfacial pressure gradient has reached its saturated value $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\lambda$ or is still building up in time as $\left.\nabla p\right|_{r=\lambda}\sim\delta p/\ell_{p}(t)$, with $\ell_{p}=\sqrt{D_{p}\,t}$, one recovers either the standard LSW coarsening law $t^{1/3}$ or the slow poroelastic coarsening law $t^{1/4}$ discussed in the main text.
+Appendix D Elastic stress boundary condition for a spherical liquid domain in an elastic, infinite and incompressible gel
+We derive the elastic stress at the boundary of a spherical solvent-rich domain of radius $\lambda$ embedded in an infinite, incompressible linear elastic gel of shear modulus $G$, when the interface is displaced radially by a small amount $\delta\lambda$.
+The interface is characterized by an isotropic surface tension $\gamma$.
+Throughout this section we emphasize an important physical point: in a gel the solvent inside the domain and the solvent in the surrounding porous network are the same continuous phase.
+Therefore, in mechanical equilibrium the hydrostatic solvent pressure must be spatially uniform.
+Only pressure gradients drive Darcy flow.
+D.1 Displacement, strain and stress
+The spherical solvent-rich domain occupies $r<\lambda$ and the elastic gel occupies $r>\lambda$. By spherical symmetry, the displacement in the gel is purely radial,
+$\mathbf{u}(\mathbf{x})=u(r)\,\hat{\mathbf{r}},\qquad r>\lambda.$ (43)
+For small strains, the only nonzero strain components are
+$\varepsilon_{rr}=\frac{du}{dr},\qquad\varepsilon_{\theta\theta}=\varepsilon_{\phi\phi}=\frac{u}{r}.$ (44)
+Incompressibility of the gel (polymer network + solvent mixture) requires
+$\mathrm{tr}\,\varepsilon=\varepsilon_{rr}+2\varepsilon_{\theta\theta}=\frac{du}{dr}+2\frac{u}{r}=0.$ (45)
+Thus
+$\frac{d}{dr}\left(r^{2}u\right)=0\quad\Rightarrow\quad u(r)=\frac{A}{r^{2}},\qquad r>\lambda.$ (46)
+For an incompressible, isotropic, linear elastic gel, the polymer stress tensor is written
+$\bm{\sigma}=-p(r)\,\mathbf{I}+2G\,\bm{\varepsilon}.$ (47)
+Here $p(r)$ is the isotropic pressure (often called the pore pressure in poroelasticity).
+Mathematically, it appears as a Lagrange multiplier enforcing incompressibility.
+Physically, it may be identified (up to an arbitrary additive constant) with the hydrostatic pressure of the solvent.
+Only its gradient has physical consequences for solvent flow through Darcy’s law.
+The pressure that enters the polymer stress tensor must equal the solvent pressure in mechanical equilibrium for an incompressible system.
+D.2 Boundary conditions and displacement field
+We impose a prescribed radial displacement at the interface:
+$u(\lambda)=\delta\lambda.$ (48)
+and determine $\delta\lambda$ below in the arrested state. For a free-standing gel, we use stress-free conditions at infinity:
+$\bm{\sigma}(r\to\infty)\to\mathbf{0}.$ (49)
+[…]
+$\delta\lambda=\frac{A}{\lambda^{2}}\quad\Rightarrow\quad A=\delta\lambda\,\lambda^{2}.$ (50)
+Hence
+$u(r)=\delta\lambda\left(\frac{\lambda}{r}\right)^{2}.$ (51)
+[…]
+ε_θθ(r) =ε_ϕϕ(r) =δλλ2r3. One immediately verifies incompressibility:
+$\varepsilon_{rr}+2\varepsilon_{\theta\theta}=0.$
+[…]
+σ_θθ(r) = -p(r) +2G δλλ2r3. Mechanical equilibrium requires that the divergence of the stress tensor vanish and in spherical coordinates this means:
+$\frac{d\sigma_{rr}}{dr}+\frac{2}{r}\left(\sigma_{rr}-\sigma_{\theta\theta}\right)=0.$ (52)
+Substitution gives
+$\frac{dp}{dr}=0,$ (53)
+so the pressure gradient is zero and we set
+$p(r)=p_{\infty},$ (54)
+a spatial constant.
+Since the deviatoric stresses vanish at infinity, the stress-free boundary condition at infinity implies $p_{\infty}=0$.
+This choice fixes a constant pressure term: only pressure differences matter physically.
+D.3 Interfacial normal stress balance and continuity of solvent pressure
+The normal stress exerted by the gel at the interface is
+$\sigma_{rr}^{\text{gel}}(\lambda)=-4G\,\frac{\delta\lambda}{\lambda},$ (55)
+while inside the solvent-rich domain the stress (negative of pressure within the solvent) is purely hydrostatic,
+$\bm{\sigma}^{(s)}=-p_{s}\,\mathbf{I}.$
+An elastic, permanently crosslinked network can coexist with solvent flow (poroelasticity). In that case, quasi-static momentum balance for the polymer skeleton includes the interphase drag:
+$\nabla\!\cdot\bm{\sigma}^{(p)}+\bm{f}_{\mathrm{drag}}=\mathbf{0},\qquad\bm{f}_{\mathrm{drag}}=\frac{\eta}{k}\,(\bm{v}_{s}-\bm{v}_{p}).$ (56)
+Therefore, “mechanical equilibrium” does not by itself imply $\nabla\!\cdot\bm{\sigma}^{(p)}=\mathbf{0}$ unless $\bm{f}_{\mathrm{drag}}=\mathbf{0}$.
+However, in the arrested elastic regime considered here, the permanently crosslinked polymer network is static in the bulk and experiences no distributed (bulk) body forces.
+Surface tension provides a typical capillary stress scale $\sim\gamma/\lambda$ (with $\lambda$ a typical radius of curvature / domain size), but it enters the problem only through interfacial traction (normal-stress) boundary conditions.
+In an elastic solid, these interfacial tractions can be balanced locally by the network’s deviatoric (non-isotropic) elastic stresses, rather than by a spatially varying pressure field of the solvent in the pores of the gel.
+Under this arrested/static assumption, the polymer bulk equilibrium (away from the boundaries) reduces to the zero net-force condition: $\nabla\!\cdot\bm{\sigma}^{(p)}=\mathbf{0}$.
+[…]
+The Young–Laplace traction condition reads
+$\sigma_{rr}^{\text{gel}}(\lambda)-\sigma_{rr}^{(\ell)}(\lambda)=\frac{2\gamma}{\lambda}.$ (57)
+Using $\sigma_{rr}^{(\ell)}=-p_{\ell}=-p_{\infty}=0$, this reduces to
+$\sigma_{rr}^{\text{gel}}(\lambda)=\frac{2\gamma}{\lambda}.$ (58)
+Substituting the elastic stress gives
+$4G\,\frac{\delta\lambda}{\lambda}=\frac{2\gamma}{\lambda},$ (59)
+which determines the displacement $\delta\lambda$
+$\delta\lambda\sim\frac{\gamma}{G}$ (60)
+(up to a numerical prefactor depending on sign convention).
+To summarize: In elastic equilibrium, the isotropic solvent pressure is spatially uniform so no Darcy flow occurs.
+The Laplace normal stress is balanced entirely by deviatoric (non-isotropic) elastic stress.
+This is fundamentally different from the viscoelastic regime of the gel, where the Laplace pressure appears as a jump in solvent pressure which then decreases as a power law within the gel.
+That solvent pressure gradient drives Darcy flow of the solvent from the solvent-rich domain into the gel, i.e., coarsening.
+In contrast, in the elastic regime, the solvent (and polymer) pressure remains uniform and the force that balances the capillary (surface tension) force is elastic and deviatoric.
+The lack of a solvent pressure gradient means that Darcy flow of the solvent, and hence coarsening is arrested.